### v2

# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

Abstract
Political actors often manipulate redistricting plans to gain electoral advantages, a process known as gerrymandering. Several states have implemented institutional reforms to address this problem, such as establishing map-drawing commissions. Estimating the impact of such reforms is challenging because each state structures bundles of rules in different ways. We model redistricting as a sequential game, where each state’s equilibrium solution summarizes multi-step institutional interactions as a single-dimensional score. We argue this score measures the leeway political actors have over the partisan lean of the final plan. Using a differences-in-differences design, we demonstrate that reforms reduce partisan bias and increase competitiveness when they constrain partisan actors. We perform a counterfactual policy analysis to estimate the effects of enacting recent reforms nationwide. Though commissions generally reduce bias, reforms that restrict partisan actors in multiple ways like removing veto points (Michigan) are much more effective than commissions where parties retain some control (Ohio).
Keywords redistricting • differences-in-differences • formal modeling • continuous treatment
1 Introduction
Democratic institutions play a crucial role in preventing political actors from pursuing policies that prioritize their own interests over the broader public good (Federalist papers 53 and 54: Madison, 1788a, ; Madison, 1788b, ). Further, increasing political polarization and the erosion of democratic norms in many countries have encouraged more recent attention to placing additional institutional constraints on policy makers (Levitsky and Ziblatt,, 2023; Little and Meng,, 2023; McCarty,, 2019). To safeguard institutions, advocates have suggested reforms that insulate democratic processes from partisan control. [^2] For example, see People Not Politicians Oregon or Ohio’s Citizens Not Politicians. For example, the Electoral Count Reform and Presidential Transition Improvement Act of 2022 represents an effort to make it more difficult for partisan actors to manipulate the presidential electoral certification process in the United States.
We study reform in American congressional redistricting, a political process often exploited by partisan actors to enact districting plans that favor their own party. This manipulation, known as partisan gerrymandering, has been widespread in the past two redistricting cycles (Kenny et al.,, 2024, 2023; Warshaw et al.,, 2022). Redistricting plans that disproportionately favor a certain party can limit how responsive a party’s share of seats in the legislature is to changes in its vote share and can reduce the electoral power of racial minorities (e.g. Canon,, 2022; Canon and Race,, 1999; Grofman and Handley,, 1991; Polsby and Popper,, 1991).
Reform efforts to limit gerrymandering are often designed to constrain partisan map drawers. They include the establishment of independent map-drawing commissions and the introduction of court oversight over proposed plans (Cain,, 2012). Estimating the causal impact of these institutional reforms, however, is challenging for three reasons. The first is the problem of treatment complexity. Redistricting reform efforts must intervene in a complex and multidimensional process. The specific rules governing each state’s redistricting process differ in many ways, including who proposes initial maps and whether or not courts can intervene. Second, states that adopt redistricting reforms may differ in both observable and unobservable ways from states that do not. This selection bias problem is common to any observational study and must be addressed to estimate credible causal effects.
Finally, the outcome of the institutional process — a redistricting plan — is also complex. For example, partisan features may be confounded by other factors such as a state’s geography and demographics (e.g., Cottrell,, 2019). If a large number of Democratic voters live in cities, any redistricting plan with compact districts may end up creating a small number of heavily Democratic-leaning districts rather than efficiently allocating Democratic votes across more districts. We must overcome this outcome complexity to accurately measure the partisan bias of each enacted plan.
In this paper, we propose a new methodological approach that addresses the above challenges to study the causal impact of redistricting reform. To deal with the treatment complexity, we first collect and standardize information about each state’s relevant laws, starting from the initial map drawer (i.e., the redistricting commission or legislature) and following through varying stalemate processes and opportunities for court intervention. We then reduce the differing districting procedures across states to a single theoretically informed parameter by modeling redistricting reform as a zero-sum sequential game. We use our original dataset of institutional procedures to characterize the players and available moves in the game. Analyzing the game allows us to measure the ability of partisan players to maximize the partisan lean of a redistricting plan. The Nash equilibrium of the game is a measure of the “leeway” that a single party has over the final redistricting plan.
We use two variations of each state’s game to produce two measures of leeway. The first is the realized leeway, which uses the observed parties of the players to compute the equilibrium. The second is the maximum leeway, which instead computes the equilibrium under one-party control. The latter allows us to measure the strength of institutions separately from party control, which varies across states and over time. We emphasize that both of these leeway variables are summary measures of our high-dimensional treatment variables of interest (i.e., institutional features) and are not functions of the outcome variable of interest. Therefore, our methodology places an observational study of institutional changes under the design-based approach of causal inference (Rubin,, 2008).
Specifically, we address the selection bias problem by using our leeway measures as a continuous treatment variable in a differences-in-differences (DiD) design applied to the 2010 and 2020 redistricting cycles. For example, states with Democratic control tend to enact more liberal policies (Caughey et al.,, 2017). This approach addresses potential confounding by comparing changes in states that have enacted reforms to those in similar states that have not, under a parallel trends assumption (Callaway et al.,, 2024). We use this model to estimate how changes in the map drawer’s leeway influence resulting plans across a set of both partisan and nonpartisan outcomes, such as the efficiency gap (Stephanopoulos and McGhee,, 2015) and responsiveness to swings in partisan preferences.
Our approach uses a formal model to summarize a multi-dimensional treatment variable, and then apply a causal inference methodology to adjust for bias due to unobserved confounding. However, we do not adopt a structural modeling approach, which would require modeling a state’s decision to adopt particular reforms for their redistricting process. The reason is that, while we are able to measure all the institutional features of redistricting processes, it is impossible to observe all the factors which affect a state’s decision to adopt redistricting reforms. Instead, we use a DiD design to address the issue of unobserved confounding.
Lastly, we deal with the issue of outcome complexity by separating the causal effect of state-specific institutional reforms from changes in political geography over time. To do this, we generate a sample of alternative redistricting plans via a simulation algorithm (McCartan and Imai,, 2023) following each state’s rules and political geography, but without regard to any partisan information (McCartan et al.,, 2022; Kenny et al.,, 2024). We use these simulated plans as a nonpartisan baseline for both 2010 and 2020 redistricting cycles, and compute the difference in outcome variables between the enacted and simulated plans. This adjusts for state-specific changes in political geography, making the parallel trends assumption more plausible.
We find that more restrictive redistricting processes reduce partisan bias by constraining map drawers. For example, changing from a single-party legislature to an independent commission leads to a reduction of about 0.5 excess seats. These effects are substantively large, as congressional gerrymandering generally results in gains of less than one or two seats per state (Kenny et al.,, 2023). Similarly, we find smaller but positive effects of constraining reforms on electoral responsiveness. We estimate that a similar change from legislature to commission would increase the share of competitive seats within a state from 25% to 38% on average.
A key advantage of our methodological approach is the ability to perform a counterfactual analysis of institutional reforms. We investigate how enacting recent procedural reforms nationwide could reduce widespread partisan gerrymandering. We quantify how the partisan bias and responsiveness of adopted plans could counterfactually change if all states nationwide adopted three kinds of commission structures currently enacted in several states: (1) an Ohio-style approach that requires supermajorities and uses a bipartisan backup commission, (2) a New York-style commission with a nonpartisan map drawer but several partisan veto points later in the process, and (3) a Michigan-style commission with a nonpartisan commission drawer, no partisan veto points, and the potential for court review.
We find that commissions can generally reduce existing pro-Republican bias, but the details of commission structure matter. In particular, unlike reforms adopted in Ohio and New York, a Michigan-style nonpartisan commission has no partisan veto points. We find that implementing Michigan-style reforms nationwide leads to an additional 22 Democratic seats, on average. Michigan-style reforms also increase electoral responsiveness whereas those adopted in New York and Ohio do not.
Beyond the substantive findings, this paper advances a new methodology for studying redistricting reform, which we hope will translate to the study of other institutional systems. Reform efforts in redistricting have produced diverse institutional changes across states. While most scholars have classified reforms into different categories (Cain,, 2012; Edwards et al.,, 2017; Warshaw et al.,, 2022; Nelson,, 2023), such an approach may miss important differences in these institutional changes.
In contrast, we use formal modeling to place these complex institutional characteristics on a continuous univariate scale to summarize how they constrain partisan actors. This theoretically driven approach, which we demonstrate accurately predicts empirical patterns, makes it possible to apply a difference-in-differences strategy to our complex setting for credible causal inference.
While our work estimates the causal effects of redistricting reform as a whole, much of existing causal analysis has focused on one part of the process, such as voter registration patterns after the Shelby decision (Komisarchik and White,, 2021) and turnout in packed districts (Fraga et al.,, 2022). Other studies have reported positive associations between the adoption of commissions and competitiveness (Carson and Crespin,, 2004; Carson et al.,, 2014; Best et al.,, 2021; Nelson,, 2023) as well as a negative association between commissions and incumbent protection (Nelson,, 2023). These studies, however, do not adjust for various confounding factors considered in our analysis.
Beyond redistricting, our methodological approach can be seen as a general strategy for the causal analysis of complex institutional reforms. By combining a game-theoretic treatment model with causal inference methods, we are able to leverage the strengths of these two approaches. A game-theoretic approach is a substantively effective way to model formal institutions with specific rules, making it possible to map multidimensional policies to a univariate summary in a theoretically-informed way. Once this summarization of various institutional reforms is done, we can apply standard causal inference methods to estimate the effects of counterfactual policies by mapping them directly to this univariate treatment variable.
Our work relates to a broader methodological literature in the social sciences. In economics, for example, Chetty, (2009) advocates the combined use of structural and reduced-form approaches via a sufficient statistic in a manner similar to our approach. In sociology, Lundberg et al., (2021) emphasize the importance of theoretically informed quantities of interest in causal analysis. Finally, in political science, Canen and Ramsay, (2023) call for the integration of rigorous theoretical and empirical approaches in causal research. We demonstrate how such an analysis can be done in the estimation of causal effects of redistricting reforms.
2 Overview of Redistricting Processes and Reforms
Before introducing our methodological framework, we provide a brief overview of redistricting processes and reforms. Every decade following the U.S. Census, states and localities use many different procedures to redraw their legislative district boundaries. For example, states differ on whether legislatures or independent redistricting commissions propose initial plans for new Congressional district maps. If these actors fail to produce a plan, state laws further vary on how these stalemates are handled. Some states pass responsibilities to a court (e.g., Virginia), a backup commission (e.g., Ohio), or a group of state party leaders (e.g., Iowa). Even plans that pass the proposal step can face a veto from other actors, such as the governor and state legislature.
Figure 1: Redistricting procedures for all 44 states with more than one district in 2020. Each vertical column indicates a separate step in the redistricting process, and nodes indicate different procedures that each state can adopt at that step. The width of each area connecting the nodes is proportional to the number of states with that specific combination of procedure at both ends. Yellow nodes indicate actors or institutions that are not explicitly partisan, while blue nodes indicate explicitly partisan actors or choices. Here, we collapse multiple potential stalemate and veto procedures into one step for visual clarity (e.g., Governor + Voters indicates the possibility of a first veto by a governor, and a second by the voters).
As explained in Section 3.1 below, we collect information about the redistricting process used in each state for the 2010 and 2020 cycles. Figure 1 summarizes this dataset and illustrates the diversity of 2020 redistricting procedures across states. For example, 29 states drew initial plans in their legislature, while 15 states used an independent redistricting commission instead. There is significant variation in the procedures following this initial draw, with seven distinct veto mechanisms across all the states. Much of this variation originates from state-specific policy campaigns that have adopted widely different goals driven by local actors (Keena et al.,, 2021). For example, California created a redistricting commission that has the power of drawing congressional district boundaries through a series of ballot propositions in the early 2000s. In recent years, several states have also implemented redistricting reforms in response to organized efforts to limit partisan influence. For example, Michigan voters approved a constitutional amendment to establish an independent redistricting commission in 2018.
These procedural differences make it difficult to attribute causal effects to particular institutional designs or reforms. Most existing work has turned to classification schemes that simplify this variation by assigning the control of map redrawing to a single drawer, typically the creator of the initial or final plan (Edwards et al.,, 2017; Carson et al.,, 2014). A more complex alternative would be to create indices that count the number of times a particular action appears in the state procedure (e.g., the number of veto points). For example, as shown in Figure 1, even states that have a commission propose their initial plan vary drastically in how the process continues afterwards.
While these simplifications make the study of complex procedures tractable, they risk oversimplifying complex procedures in two ways. First, attributing institutional outcomes to a single actor ignores the fact that the final redistricting plan is produced through a series of steps. Take, for example, the New York redistricting process in 2020. An independent bipartisan commission had the power to propose an initial plan, but failed to agree on a single plan. State law required the stalemated process to move to the legislature, which adopted a plan that Republican and civil rights groups criticized as a Democratic gerrymander. These groups challenged the adopted plan in a series of lawsuits, and in 2022 the New York Court of Appeals struck down the plan and tasked a court-appointed special master with drawing a remedial plan. Though the final plan scores well on quantitative fairness metrics (Kenny et al.,, 2023), simply classifying a “court” as the sole map drawer for New York in 2020 ignores the partisan interests involved, failing to capture the complexity of the procedures that led to the final plan.
Second, classification schemes can overlook strategic interactions, where the behavior of certain actors depends on the presence or characteristics of others. For example, a commission might draw a different map if it knew that the map could later be reviewed by a court. Or, the potential of a governor’s veto may limit the likelihood of a partisan gerrymander by the state legislature, but not in cases where the governor and legislative majority share a partisan affiliation. More detailed coding of procedural schemes can account for some of these interaction effects and increase realism, but will necessarily decrease statistical power, making it more difficult to estimate causal effects.
Thus, we face a methodological dilemma: while some simplification of institutional features is necessary, common approaches to doing so obscure critical characteristics of redistricting processes. In the next section, we propose a theoretically grounded approach that models redistricting as a zero-sum sequential game and uses its Nash equilibrium as a treatment variable to summarize these complex institutions.
3 A Theoretical Model of Institutional Leeway
To study the impact of redistricting reforms and processes, we first develop a standardized set of 14 institutional features that capture the most important actors in congressional redistricting processes across states. We collect data on these features so that it is possible to compare redistricting processes consistently across states, despite the differences in the laws and bodies governing redistricting in each state.
Next, we develop a sequential redistricting game to summarize these features in a theoretically informed manner. Specifically, we use the Nash equilibrium of this game as a one-dimensional measure of the institutional leeway political actors have over the partisan lean of the resulting redistricting plan. Finally, we empirically validate the proposed measure by demonstrating that it is not particularly sensitive to model specification and predicts redistricting outcomes well. We emphasize that the proposed measure is a theoretically motivated summary of our high-dimensional treatment variable of interest. Since this measure does not use the outcome variable, we are able to apply the standard causal inference methodology to adjust for observed and unobserved confounding factors.
3.1 Data on the relevant institutional features
Figure 2: Prototypical game tree used to model redistricting in every state. States differ in which party, if any, controls each node, and which nodes are present in the state’s process.
Our standardized coding of relevant institutional features is based on the prototypical redistricting process shown in Figure 2. First, an initial map drawer proposes a plan that may be vetoed by other actors. If the plan is not vetoed, it can be challenged in court or not. If it is vetoed, there is another round of map drawing. If a plan is vetoed twice, or if the initial map drawer cannot agree on a plan, then a different institution (often a court) must resolve the stalemate and adopt a plan.
Every state’s process can be described as a subset of this prototype. For example, in Michigan, a commission draws congressional districts, resolves any stalemates, and there is an explicit mechanism for state court review. Thus, Michigan’s process would be described by the “Drawer” step in Round 1 and “Stalemates” steps only in Figure 2.
For each state, we record which institutional body, if any, acts at each step and which party, if any, controls that institution. We also collected additional information relevant to modeling redistricting processes and court review, such as whether the state’s redistricting plans were subject to DOJ preclearance before 2013, [^3] The Shelby decision was handed down in 2013, ending DOJ preclearance. and which institutional body ended up drawing the plan that was used in the first postcensal elections. Most procedural details are straightforward, and the information for coding is readily available in public data from each state (see e.g., https://redistricting.lls.edu/). Furthermore, many states have similar redistricting procedures that are easily classified under the categories we defined earlier.
[…]
3.2 The redistricting game
Coding the details of each state’s redistricting process preserves important procedural information, compared to categorizing each state into a small number of groups such as “legislature-controlled” and “independent commission.” However, the detailed coding presents a challenge for causal inference, since the treatment—a state’s redistricting process—is now high-dimensional. Out of the 87 state-decade processes we code, [^4] Only 43 states had two or more congressional districts in the 2010 cycle; for 2020, Montana did as well, bringing the total to 44. there are 58 distinct combinations of procedural variables, 41 of which are completely unique.
The combination of high-dimensional treatment and a limited sample size means that there may not be enough information to estimate the causal effect of changing from one specific configuration of procedural variables to another, without further assumptions.
We address this methodological challenge by leveraging two basic substantive assumptions about the redistricting process. First, each party aims to draw a map that favors it as much as possible, and second, the parties are constrained by statutory and constitutional rules in doing so. Specifically, we treat the redistricting process depicted in Figure 2 as a sequential zero-sum game [^5] The zero-sum assumption rules out cooperation between the parties to, e.g., protect incumbents. Given the role of the game here as a summary of institutional features rather than a structural model that predicts the outcome, we believe this is an acceptable limitation. with two players—the Democratic and Republican parties—each trying to maximize the degree to which the drawn plan favors their party.
In each state, the nodes in the game tree can be controlled by different parties, or by neither party (e.g., when a supermajority is required to adopt a plan and neither party has supermajority control). Any split-control nodes, as well as a node for state court action, are considered moves by nature. Moves at one node do not affect the set of actions available at other nodes. Some nodes involve discrete choices such as whether to veto a plan or not, others are labeled “choose bias,” meaning that the player at that node draws a plan with a chosen amount of partisan bias favoring either party. This bias is exactly the utility for the party of the chosen plan: a plan with bias $x$ is worth $x$ to the Republicans and $-x$ to the Democrats. We need not quantify exactly what the utility measure is as a function of a specific plan chosen; it suffices to let the parties try to maximize an abstract univariate measure of partisan bias. We let the bias score range from $-4$, indicating a maximum Democratic advantage, to $+4$, indicating a maximum Republican advantage.
For concreteness, consider the case of redistricting in Michigan, as described earlier. The first move belongs to the independent commission. It has to pick the amount of partisan bias $-4<x<4$ in the plan that it adopts. If this plan is ultimately the final enacted plan, Republicans receive utility $x$ and Democrats receive utility $-x$. The stalemate move also belongs to the commission.
If a plan is adopted in the first round, it proceeds to a possible court review. The court may decide to accept a legal challenge, decide in favor of the plaintiffs, and redraw the map; this choice is considered a move by nature. If the court review results in a redrawn map with partisan bias $x^{\prime}$, then the Republicans receive utility $x^{\prime}$ and the Democrats receive utility $-x^{\prime}$. In Michigan, court review is explicitly allowed on partisan grounds and challenges under the federal Voting Rights Act are possible, so there is moderate probability that the commission-adopted plan will be overturned. If no plan is enacted by the commission in the first round, the commission is again tasked with drawing a plan. If the commission fails to enact a plan again, courts must step in and redraw district lines to ensure compliance with federal constitutional “one person, one vote” apportionment requirements. This is also considered a move by nature.
[…]
All in all, the game specification depends on 19 parameters which govern the moves by nature, with most of these parameters relating to the court challenge process. Rather than fix these parameters to arbitrary constants, we place a prior distribution on each parameter over a range of probable values. We simulate 100 different draws from this joint prior distribution; each draw generates a slightly different game specification. We then average our results across the random draws. [^6] We average, rather than propagating each draw through the estimation step. This has computational benefits and reduces the complexity of an outcome model (the outcome model would need to account for a different scale of treatment values across different draws). More importantly, however, the variation in the estimated equilibria, as described below, is quite small compared to other sources of uncertainty.
3.3 Equilibrium solution as a treatment variable
To effectively summarize a high-dimensional treatment, we use the subgame perfect Nash equilibrium of the redistricting game. The utility of the equilibrium solution captures the expected partisan bias of plans that arise out of a state’s redistricting process, under the current party control of the state’s institutions. All of the multi-step institutional interactions and negotiations that might happen as part of the redistricting process are thus reduced to a univariate score. The upshot is that the 14 procedural variables can be reduced into a single variable that measures the leeway political actors have over the partisan lean of the final redistricting plan.
[…]
The expected outcome of both the stalemate-then-court and court-alone choices are determined by the parameters governing the moves by nature. The overall equilibrium is a bias of $-0.03$, indicating that there is essentially no bias in favor of either party. This solution process is automated across all of the states and is carried out on each of the 100 different draws from the prior on the game’s parameters.
For our causal analysis, the first treatment measure is exactly the game’s Nash equilibrium, averaged across the prior. This measure depends on which parties control each node in the game tree. We also generate a second treatment measure that does not depend on the current party control. This is calculated by assigning a single party (here, the Democrats [^7] The game is completely symmetric between parties, with one exception: due to racial polarization in the U.S., challenges to redistricting plans under the VRA are much more likely for Republican-leaning plans, and when these challenges prevail the remedy almost always results in a plan more favorable to Democrats. In keeping with the goal of calculating the maximum leeway in each state, we therefore assign states to Democratic control under the counterfactual scenario here. Assigning to Republican control instead would slightly reduce the calculated leeway for those states previously subject to DOJ preclearance.) to control each node in the game tree that belongs to a partisan actor—legislature, governor, or partisan commission—and then recalculating the average Nash equilibrium. We refer to this equilibrium as the maximum leeway of a state’s redistricting process, since it captures the expected bias under a worst-case partisan outcome where all of the levers of state government are controlled by one party. The first treatment measure we refer to as the realized leeway, since it depends on the realized values of the party control for each state institution.
Figure 3: Treatment values for each state in 2010, with values for 2020 indicated by arrows, where different. States in orange are those which experienced a reform to their redistricting procedures, either by legislation, constitutional amendment, or a court ruling that allowed for state court review of alleged partisan gerrymanders.
Figure 3 visualizes these treatment measures for all the states we study; an arrow indicates a change in the treatment value from 2010 to 2020. The realized leeway scores (game equilibria) cover the entire range of possible partisan biases, from West Virginia in 2010 with complete Democratic control of state government, to Wisconsin, Iowa, Kansas, Utah, and Nebraska, where Republican trifectas were unconstrained by Democrats or by the VRA in both cycles. The maximum leeway scores likewise span the range of possible bias, but take on fewer values, since the various actual combinations of partisan control of state institutions are no longer considered.
As might be expected, states with independent redistricting commissions, such as the running example of Michigan, have a realized and maximum leeway of zero. In 2010, Michigan redistricting was entirely controlled by the legislature and governor, who were both Republican. By 2020, the process was reformed to have a strong independent commission. In the new process, the commission is bound by strict criteria, the process allows for court review, and even the commissioners themselves are selected in part by a lottery system. With the reform, Michigan’s maximum leeway changed from the highest observed value to zero, meaning that partisan actors have no control over redistricting outcomes, on average.
Other states’ scores are similarly interpretable, where introducing a commission (e.g., New York, Virginia, or Colorado) drastically reduces its leeway to match states with similar sets of rules in 2020. States that had intervening litigation to clarify the interpretation of state redistricting rules (e.g., New Mexico, North Carolina, or Pennsylvania) see appropriate changes in leeway, even absent a commission. Further, in our party-signed treatment, states that had a flip in party control of the whole system see large changes in leeway in the correct direction (e.g. West Virginia). Finally, states with minor changes to the total control of state institutions see minor changes in leeway (e.g., North Carolina).
3.4 Validation
[…]
The model was designed to be flexible enough to capture the actual redistricting processes in each state. Any misspecification is therefore due to how the moves by nature are specified, or in the underlying setup of two opposing partisan actors competing on a single zero-sum dimension. First, we find that the treatment values are not sensitive to specific parameter values, up to monotonic transformations. Across the 100 random draws from the prior, the average pairwise Spearman correlation between the Nash equilibria for each state is above 0.99.888i.e., for each pair of random draws, we calculate the Spearman correlation between the two vectors of calculated state equilibria, then average these correlations across all pairings. This also gives us confidence that the specific choice of the prior is not influencing the results.
Figure 4: Measures of partisan advantage versus treatment values for states’ enacted plans for the 2010 and 2020 redistricting cycles. Points are slightly jittered to avoid overplotting.
Second, as shown in Figure 4, there is a substantial correlation in the expected direction between the treatment values and the measures of partisan advantage. The left panel plots the expected share of seats won by Republicans versus the realized leeway measure; recall that the leeway measure is positive for plans that favor Republicans. The right panel shows the partisan bias (King and Browning,, 1987), measured at the state’s baseline vote share, versus the realized leeway. The greater values of partisan bias correspond to plans that systematically favor one party, with positive values favoring Republicans.
For both outcome measures, the purely a priori model predictions of the expected partisan bias do in fact correlate with the actual partisan bias of the plans that come out of each state’s redistricting process. To what extent these correlations can tell us about the causal effects of reforms is of course another matter, and the primary question addressed in the next section.
Table 1: Correspondence between the institution that drew the final redistricting plan for each state and the most likely outcome based on the equilibrium path of the redistricting game.
Most likely in equilibrium
Final drawer Legislature Commission Court Total
Legislature 31.9 0.0 19.1 51
Commission 0.0 18.9 3.0 22
Court 1.7 0.0 12.3 14
Total 33.6 18.9 34.4 87
We also compare the equilibrium path in the model with the actual institution that drew the final map in each state. Because the moves by nature in the game are random, the actual equilibrium path may be a probabilistic mixture over multiple possible paths. We average these probabilities over the 100 draws of the game to arrive at an overall probability that a legislature, commission, or court draws the final map in each state, given its procedure and party control.
Table 1 compares the most likely final map-drawer, according to these probabilities, to the actual institution that drew the final redistricting plan. The correspondence is generally excellent, considering that the predictions are made based on theory alone. There is a tendency, however, for the model to predict court-drawn maps when they are in fact drawn by legislatures. The overall likelihood of a court intervention is itself a model parameter that is varied across random draws, and as noted above the ranking of the states by leeway is basically unchanged by different parameter values. Therefore, we expect this tendency to over-predict court intervention to at most slightly understate the equilibrium value for states with court review, since a higher likelihood of intervention would tend to pull the equilibrium towards zero.
Finally, it is natural to wonder whether a more accurate model could be obtained by estimating the model parameters from the observed data rather than specifying their prior distributions. That is, one could find the parameter values that maximize the correlation between the treatment and the outcome. We do not take this approach to cleanly separate the estimation of causal effects from the construction of a treatment variable (Rubin,, 2008). This enables us to employ a causal identification strategy that does not assume the correct specification of the game-theoretic model. Instead, we use the model to summarize a high-dimensional treatment variable in a theoretically informed way without looking at the outcome variable. [^9] We did attempt to fit the game parameters to the observed data for the 2010 cycle alone via maximum likelihood, but we found that the observed data were only minimally informative about most of the parameter values. This is good news, insofar as it implies that a wide range of parameter values are all compatible with the observed data.
4 Estimating the Causal Effects of Institutional Leeway
4.1 Difference-in-differences design
To estimate the causal effects of changes in leeway on redistricting outcomes, we use a differences-in-differences (DiD) design with a continuous treatment variable. This strategy addresses potential confounding by comparing changes in states that have enacted reforms to changes in similar states that have not modified their redistricting process across the 2010 and 2020 redistricting cycles. We assume that, in the absence of reforms, the states with institutional changes would have experienced the same trend in outcome variables as those states without reforms.
Formally, let $\mathbf{Z}_{it}$ be the 14-dimensional vector of institutional features for state $i$ at time $t$ discussed in Section 3.1. We use $t=0$ and $t=1$ to denote 2010 and 2020 redistricting cycles, respectively. Let $u^{*}(\mathbf{z})$ represent the equilibrium outcome (game utility) for the average Nash equilibrium of the redistricting game described by the process $\mathbf{z}$. Then, we can define a univariate treatment variable $D_{it}=u^{*}(\mathbf{Z}_{it})$ for each state-decade, which represents the “dose” of institutional leeway given to political actors.
The key assumption made by our use of the Nash equilibria as the treatment variable is that the institutional features affect the outcome only through this treatment variable, i.e.,
$Y_{it}(\mathbf{z})=Y_{it}(\mathbf{z}^{\prime})\mbox{\quad for any\quad}\mathbf{z},\mathbf{z}^{\prime}\text{ with }u^{*}(\mathbf{z})=u^{*}(\mathbf{z}^{\prime}),$
[…]
Our target estimand is the conditional average treatment effect (CATE) for a change from the treatment level $d$ to $d^{\prime}$ given a set of covariates $\mathbf{X}_{i}$ that are not affected by the treatment. The CATE is defined as,
$\mathrm{CATE}_{\mathbf{x}}(d^{\prime},d)=\operatorname{\mathbb{E}}[Y_{i1}(d^{\prime})-Y_{i1}(d)\mid\mathbf{X}_{i}=\mathbf{x}].$
Estimating $\mathrm{CATE}_{\mathbf{x}}(d^{\prime},d)$ for any pair of values $d^{\prime},d$ and covariate values $\mathbf{x}$ requires the identification of a full dose-response curve. Our identification approach relies on a strong conditional parallel trends assumption introduced by Callaway et al., (2024). Specifically, we assume that the average change in outcomes for states experiencing a change in treatment dosage from $d$ to $d^{\prime}$ depends only on the observed covariates, and not the observed dosage value.
Our covariates include a range of variables that existing work in the redistricting literature identifies as predicting redistricting outcomes and are not affected by redistricting reforms. We include (pre-treatment) 2008 Democratic presidential vote share, an indicator for being in the South, [^11] Southern states are defined as Alabama, Arkansas, Florida, Georgia, Kentucky, Louisiana, Mississippi, North Carolina, South Carolina, Tennessee, Texas, Virginia, and West Virginia. This set corresponds to Confederate states, plus Kentucky and West Virginia, which also had a history of Democratic dominance in the 20th century and notable differences between party preferences at the state and national levels. the logarithm of the number of districts in 2020, the change in the number of districts between 2010 and 2020, the logarithm of the average number of state corruption convictions by year between 2000 and 2010, and an indicator for whether states allow ballot initiatives. For the maximum leeway treatment, which uses no information about partisan control of state institutions, we also control for changes in partisan control of the map-drawing body and of the state supreme court.
Formally, we require the following conditional parallel trends assumption for all $d,d^{\prime}$:
$\operatorname{\mathbb{E}}[Y_{i1}(d^{\prime})-Y_{i0}(d)\mid\mathbf{X}_{i}=\mathbf{x}]=\operatorname{\mathbb{E}}[Y_{i1}(d^{\prime})-Y_{i0}(d)\mid\mathbf{X}_{i}=\mathbf{x},D_{i0}=d,D_{i1}=d^{\prime}].$
This assumption is analogous to the parallel trends assumption in the traditional binary difference-in-differences design, but differs in that it refers to changes in dosage across a continuous measure, rather than a single level change from zero. This conditional parallel trends assumption allows us to identify our estimand as
$\mathrm{CATE}_{x}(d,d^{\prime})=\operatorname{\mathbb{E}}[Y_{i1}-Y_{i0}\mid\mathbf{X}_{i}=\mathbf{x},D_{i0}=d,D_{i1}=d^{\prime}]-\operatorname{\mathbb{E}}[Y_{i1}-Y_{i0}\mid\mathbf{X}_{i}=\mathbf{x},D_{i0}=d,D_{i1}=d],$
where $Y_{it}=Y_{it}(D_{it})$ is the observed outcome. Averaging over the marginal distribution of $\mathbf{X}_{i}$, we can also estimate the average treatment effect (ATE).
4.2 Adjusting for political geography
To further increase the credibility of the conditional parallel trends assumption above, we adjust for the changes in political geography between the two redistricting cycles. Partisan preferences and the geographic distribution of voters can change in different ways within each state over time. For example, Michigan, which experienced reform, became more Republican during our sample period, while Georgia, which did not, became more Democratic. Without this adjustment, an estimated effect of the Michigan reform would be biased by the differential change in the states’ political geographies.
To make the adjustment, we use representative sets of nonpartisan redistricting plans within each state that respect each state’s specific redistricting rules. These simulation samples were separately generated for 2020 (McCartan et al.,, 2022) and 2010 (Kenny et al.,, 2024) using the algorithm of McCartan and Imai, (2023). We subtract the mean outcome in each state’s simulated sample, denoted by $\widetilde{Y}_{it}$, from the observed outcome $Y_{it}$ before estimating the causal effects under the DiD design. We argue that this subtraction of the outcomes based on simulated baseline plans from the observed outcomes accounts for the change in political geography in each state.
One limitation of basing simulations off of each state’s rules is that any changes in outcome variables from 2010 to 2020 that are due to changes in rules would be subtracted away. Since redistricting rules are sometimes changed as part of larger reform efforts, part of the overall effect of redistricting reform would not be included in the measured effect, attenuating the overall estimated effects. However, these rules changes occurred in only a few states, and their effect on partisan outcomes was measured by Kenny et al., (2023) and found to be minimal.
Thus, we assume that once we adjust for the state-specific change in political geography, states with different changes in institutional features have parallel trends in the potential outcomes. Formally, if we denote the difference between the potential outcome and the simulated outcome as $\Delta Y_{it}(d)=Y_{it}(d)-\widetilde{Y}_{it}$, our conditional parallel trends assumption becomes
$\operatorname{\mathbb{E}}[\Delta Y_{i1}(d^{\prime})-\Delta Y_{i0}(d)\mid\mathbf{X}_{i}=\mathbf{x}]=\operatorname{\mathbb{E}}[\Delta Y_{i1}(d^{\prime})-\Delta Y_{i0}(d)\mid\mathbf{X}_{i}=\mathbf{x},D_{i0}=d,D_{i1}=d^{\prime}]$
for all $d,d^{\prime}$. Then, as above, the CATE is identified as
$\mathrm{CATE}_{\mathbf{x}}(d,d^{\prime})=\operatorname{\mathbb{E}}[\Delta Y_{i1}-\Delta Y_{i0}\mid\mathbf{X}_{i}=\mathbf{x},D_{i0}=d,D_{i1}=d^{\prime}]-\operatorname{\mathbb{E}}[\Delta Y_{i1}-\Delta Y_{i0}\mid\mathbf{X}_{i}=\mathbf{x},D_{i0}=d,D_{i1}=d],$
where $\Delta Y_{it}=Y_{it}-\widetilde{Y}_{it}$.
4.3 Estimation
[…]
In terms of the prior specification, the intercept receives a weakly informative prior of $\mathcal{N}\quantity(0,(2.5\cdot\sigma_{y})^{2})$, where $\sigma_{y}^{2}$ is the variance of $\Delta Y_{t}$. Each main effect coefficient receives a weakly informative prior of $\mathcal{N}\quantity(0,(0.75\cdot\sigma_{y}/\sigma_{x})^{2})$, where $\sigma_{x}^{2}$ is the variance of each covariate $x$. The prior standard deviation is further scaled by 0.25 for interaction terms, following the discussion in Gelman et al., (2008). Finally, the residual standard deviation receives an $\operatorname{Expo}(1/\sigma_{y})$ prior. Computation is carried out via NUTS-HMC as implemented in the Stan probabilistic programming language (Stan Development Team,, 2024).
5 Estimated Causal Effects of Redistricting Reforms
The outcome of the redistricting process is a complete congressional districting plan. A plan can be evaluated in a number of ways—how many seats each party is expected to win, how many of the seats are competitive, how well the partisan composition of the delegation matches the voters’ preferences, and so on. We begin this section by introducing quantitative measures of several partisan and nonpartisan aspects of districting plans. These constitute the outcome variables for the causal estimates we present in the second half of the section.
5.1 Outcome measures
Our measures can be divided into two buckets: nonpartisan outcomes that quantify how much gerrymandering is present or how responsive a districting plan is to shifts in public opinion, and partisan outcomes that capture which party is advantaged by a districting plan. For the nonpartisan outcomes, we also use the absolute value of the realized leeway as an additional treatment variable so that neither treatment nor outcome considers the partisan direction.
All of these measures are calculated from district-level election results. Since there is significant exogenous variation in election results due to swings in the national political environment, rather than using actual 2012 and 2022 House election results to evaluate districting plans, we use a statistical model to estimate the distribution of election results across future hypothetical elections. A statistical election model also allows us to apply these same measures to the simulated districting plans, under which no elections have taken place.
We adopt the model from Kenny et al., (2023), which assumes district election outcomes can be decomposed into a baseline district-level vote share plus district-specific and national swings. [^12] The full details of this approach can be found in the Methods section and SI Appendix, Section B of Kenny et al., (2023) The model is closely related to the stochastic uniform partisan swing model of Gelman and King, (1994) and the congressional model of Ebanks et al., (2023). As the baseline district-level vote, we use the 2008 presidential results for the 2010 redistricting cycle and an average of the 2016 and 2020 presidential results for the 2020 cycle, each logit-shifted so that the national partisan vote is exactly 50/50 (Voting and Election Science Team,, 2018, 2020). As in Kenny et al., (2023), we use historical congressional election data since 1976 to estimate the variance of district and national swings. We calculate exact expectations of all of our outcome measures against the predictive distribution of the model via numerical integration.
Our primary partisan measure of districting plans is the expected number of seats won by the Republican party. As discussed in Section 4, we subtract the average outcomes in the simulated baseline redistricting plans (McCartan et al.,, 2022; Kenny et al.,, 2024) from the observed outcome in the enacted plan to adjust for the state-specific change in political geography, which is a potential confounding factor (Cottrell,, 2019). After this adjustment, the seat outcome ranges from $-1.57$ in Illinois in 2020 (favoring Democrats) to $1.78$ in Texas in 2020 (favoring Republicans). These simulation-differenced seats can be directly interpreted as a measurement of bias due to partisan gerrymandering, with positive values indicating a Republican bias beyond what would be expected based on the state’s political geography alone.
The expected number of Republican seats is an interpretable measure, but may not be completely comparable across states. Depending on political geography and especially the total number of districts in each state, the natural variation in the number of seats outcome may vary significantly between states. To address this, we also include as an outcome measure the simulation $z$-score of Republican seats, which is calculated by taking the simulation-differenced seats outcome and dividing it by the standard deviation of Republican seats in the simulation set (Kenny et al.,, 2023). This puts all the states’ outcomes on a common scale, increasing the plausibility of the parallel trends assumption and the homoskedasticity assumption of the estimation model.
For a nonpartisan outcome measure, we take the absolute value of these two partisan measures. Both the absolute difference and the absolute $z$-score of Republican seats measure how far the enacted plan deviates from the nonpartisan simulation baseline in terms of partisan composition. [^13] For the transformed outcomes based on $z$-score and absolute value, we do not interpret the estimated causal effects in terms of the original outcome variable (i.e., the expected number of Republican seats). Instead, the estimates should be interpreted as the effects on the transformed outcomes. This also means that the parallel trends assumption is assumed to hold on the transformed outcome variable.
We also study two more outcome measures based on the efficiency gap, which is a popular measure of partisan bias calculated as the difference between the wasted votes of the two parties (Stephanopoulos and McGhee,, 2015). Wasted votes are votes cast for losing candidates and votes cast for winning candidates beyond the threshold needed to win. The idea behind the efficiency gap is that a fair plan should have similar numbers of wasted votes for both parties. The partisan-signed version of the efficiency gap indicates which party benefits from the plan in terms of vote efficiency; we use positive values to indicate pro-Republican bias. Unlike with Republican seat share, we do not subtract the average simulated value from the partisan-signed efficiency gap measure. This is because the efficiency gap is already designed to have a universal scale, where 0 indicates partisan fairness; subtracting the simulated average would change the definition of a fair baseline for this measure. To create a nonpartisan measure, as with Republican seats, we take the absolute value of the difference between the enacted and average simulated plan’s efficiency gap. Smaller values indicate a fairer plan by this measure (closer to the normatively preferred value of 0).
Finally, we also measure the responsiveness of districting plans to changes in the national electoral environment. Responsiveness is measured as the rate of change in the share of Republican seats given an infinitesimal change in Republican vote share nationwide. Responsiveness is closely linked to the presence of competitive seats; the more competitive seats there are, the larger the change in seat share will be for a given shift in vote share. Indeed, electoral responsiveness is the primary motivation for the creation of competitive congressional districts.
The responsiveness measure ranges from 0.044 in Wyoming in 2020 to 7.91 in New Hampshire in 2020, with most plans’ values lying between 0.5 and 3. The interpretation of these values is as follows; in New Hampshire, a 1pp increase in a party’s vote share leads to a 7.91pp change in the party’s expected seat share. This relatively large increase makes sense in context—both of the state’s congressional districts have Republican vote share within a few points of 50%. As with the other measures, we subtract the mean responsiveness of the simulated plans from the enacted plan’s responsiveness when estimating causal effects.
[…]
5.2 Findings
[…]
Figure 5: Fitted model coefficient estimates for the Republican seat outcome measure (left) using the maximum leeway treatment, and estimated conditional average treatment effects for each reformed state’s covariate combination plotted against the state’s dose (right). The model-based dose-response curve is underlaid in blue. 80% and 95% credible intervals are shown throughout.
Given the fitted outcome model, we can estimate the conditional average treatment effect (CATE) for any given starting and final value of maximum leeway and any combination of covariates. The right panel of Figure 5 plots these CATEs for each state that experienced a change in maximum leeway between 2010 and 2020, using the specific covariates of that state. There is a clear dose-response relationship between the decrease in maximum leeway and the decrease in the expected number of Republican seats. [^14] This pattern was also observed in robustness checks that used a flexible Bayesian Additive Regression Trees (BART) model for estimation. A state with Michigan’s dose (nearly 4) has an estimated CATE of around $-1.25$, with 95% credible interval (–2.0, –0.4), and so would be expected to gain just over 1 Democratic seat, on average. Similarly, a state with Ohio’s dose (just over 1) and covariates would be expected to gain about half of a Democratic seat (–0.6, –0.1). Motivated by this pattern, we also include in the right figure a dose-response curve calculated by calculating a CATE for dose level and averaging over the observed covariate distribution. [^15] Because of the interaction between the dose variable and the covariates (which do not have mean zero), the overall slope estimate and its variance differ from the estimate and variance of the coefficient on Dose in the linear model.
Since the dose-response pattern is well approximated by a linear relationship, we can summarize it by its slope, a quantity also known as the average causal response (ACR). The product of the ACR and a given dose corresponds to the estimated causal effect of that dose, averaged across states.
Figure 6 presents the ACR for Republican seats and the other outcome variables, using both the maximum and realized leeway treatment values. As mentioned above, for the nonpartisan outcomes, we also use the absolute value of the realized leeway as a treatment variable. The right side of the figure plots the ACR estimates in terms of the standard deviation of the respective outcome variable per unit change in treatment. This means that a point estimate of 0.5 is interpreted as an increase of 0.5 standard deviations of the outcome variable for a 1-unit change, or an increase of 2 standard deviations for a 4-unit change.
Figure 6: Average causal response (ACR) of leeway on redistricting outcomes. The points correspond to the mean estimated ACR, while the lines represent 80% and 95% credible intervals. Intervals are colored by the treatment variable used. The numbers in the columns display the mean ACR on each outcome’s response scale. The estimates and intervals on the right are displayed in units of outcome standard deviations, to allow for comparability between outcomes. For partisan outcomes, a positive number indicates a pro-Republican effect and a negative number indicates a pro-Democratic effect for a positive dose.
Across the three nonpartisan outcomes measuring the amount of gerrymandering, we find a consistent effect of leeway. The estimated ACRs are all positive, meaning that an increase in leeway leads to a greater gap between the enacted plan and the simulated baseline. That is, states that reduce leeway through reform efforts are expected to rein in partisan gerrymandering. The magnitude of the effect is substantial: a reform effort like Michigan’s, with a decrease in the maximum leeway of 4 units, would be expected to reduce the difference in Republican seats from the baseline by about 0.5 seats (–0.006, 1.040), or 1.7 simulation standard deviations (–2, 5)—enough to move an extreme partisan gerrymander to a fair plan. The ACRs for the absolute realized leeway are very similar to those for the maximum leeway; the ACRs for the fully signed realized leeway are positive but smaller in magnitude, and have enough uncertainty that the 80% credible intervals cross zero. Thus, there is some weak evidence that giving Democrats more control of the redistricting process by increasing their leeway or reducing Republican leeway would also reduce gerrymandering.
We also find that reforms increase electoral responsiveness. The point estimates of the ACRs for both the maximum leeway and the realized leeway agree; the posterior probability that the effect is negative is 97% and 96%, respectively. The estimated effect for the absolute realized leeway is near zero. We estimate that a reduction in the maximum or realized leeway of 4 units would increase responsiveness by 1.1 (–0.042, 2.142), meaning that a 1pp increase in a party’s vote share would translate to an additional 0.6pp in seat share on top of the average 2pp increase in seat share across all states.
Another way to interpret the effect on responsiveness is in terms of the share of competitive seats, where a seat that is counted as competitive in proportion to how close its baseline partisanship is to 50%. This is equivalent to a linear rescaling of the responsiveness outcome. [^16] Specifically, we measure the competitiveness of a single seat as the derivative of the election model’s win probability for that seat at its baseline vote, divided by the maximum derivative, which is obtained at a 50/50 baseline vote. Based on our election model, a district with 50/50 baseline vote is counted as 1.0 competitive seats, and a district with 60/40 baseline vote is counted as just 0.14 competitive seats. This approach avoids arbitrary cutoffs in what counts as a competitive district. In terms of competitive seats, we estimate that a reduction in leeway of 4 units would increase the share of competitive seats in an average state from 25% to 38% (24%, 51%).
In terms of partisan outcomes, the estimated ACRs on Republican seats are all positive, for both realized and maximum leeway. This is in line with an interpretation that increasing leeway benefits the party in control, as one might expect. Recall that the doses from Figure 3, for the realized leeway estimates, a dose of 4 represents approximately going from a Democratic legislature to an independent commission or an independent commission to a Republican legislature. For such a dose of 4, we would expect Republicans to gain around 0.7 seats (0.21, 1.18), or 7.2 simulation standard deviations (2.4, 12.2)—again, a substantial effect. We do not find a consistent effect on the signed efficiency gap outcome, which was not adjusted using the simulations.
6 Redistricting Commissions: Counterfactual Policy Analysis
What if redistricting reforms were adopted nationwide to constrain partisan actors? Redistricting commissions are the most commonly adopted map-drawing reform. A total of 15 states, including Michigan, New Jersey, and Arizona, currently use some form of commission. Advocates often argue that commissions can lead to fairer redistricting plans by further removing legislators from the process (e.g., American Academy of Arts & Sciences Commission on the Practice of Democratic Citizenship,, 2020). [^17] See also People Not Politicians Oregon or Ohio’s Citizens Not Politicians
In this section, we conduct a counterfactual policy analysis of the nationwide adoption of redistricting commissions. Though hypothetical, counterfactual policy analyses are useful tools for evaluating institutional reform proposals (Cervas and Grofman,, 2019). Our approach is well suited for this analysis, in part because commissions can vary drastically in the extent to which they limit the influence of partisan actors. We use our causal models to predict electoral outcomes under a series of hypothetical scenarios, in which every state adopts a redistricting commission with particular institutional powers. Although each of the 15 states uses a redistricting commission with different structures and rules, our model is able to characterize these different commissions in terms of their partisan leeway and use these treatment values to predict electoral outcomes under a given commission structure.
We study three types of commission structures currently enacted in several states: (1) a New York-style commission with a nonpartisan map drawer and several partisan veto points; (2) an Ohio-style reform with legislature-drawn map and several partisan and bipartisan veto points and stalemate procedures, including a bipartisan commission stalemate procedure; (3) and a Michigan-style reform, with a nonpartisan commission, no partisan veto points, and the potential for court review. [^18] Although states could change their redistricting processes in the future, we use shorthands like “Michigan-style” to refer to the current (as of 2024) institutional design of each state’s commission structure. Though each of these states has a commission of some kind, they differ in important ways — for example, Michigan’s elimination of partisan veto points removes the ability of partisan actors to veto unfavorable plans.
Figure 7: Commissions Can Reduce Partisan Bias. The figures show three predicted seats-votes curves if all US states adopted new redistricting institutions with: (1) a New York-style commission with a nonpartisan map drawer and several partisan veto points; (2) an Ohio-style legislature-drawn map and several partisan and bipartisan veto points; and (3) a Michigan-style reform, with a nonpartisan commission, no partisan veto points, and the potential for court review. Hypothetical commission structures are plotted as orange lines (with 80% and 95% credible intervals), with reference lines for actual plans for both 2020 and 2010 in black.
We visualize the impact of each reform structure using a seats-votes curve. Seats-votes curves are often used to study a measure of partisan bias, operationalized as the difference in the share of seats and votes at a given point on the curve (Tufte,, 1973; King and Browning,, 1987; Katz et al.,, 2020). A seats-votes curve for which a 50% vote share translates exactly to a 50% seat share represents the baseline for partisan bias. We approximate the seats-votes curve around 50% vote share as a line, and estimate the effect of reforms on its slope and intercept separately. Up to a linear transformation, the slope corresponds to responsiveness and the intercept to seat share. These two outcome variables are differenced using the simulation sets, as discussed in Section 4.
We find that nationwide adoption of commissions would generally reduce partisan bias, and that reforms that place additional restrictions on partisan actors (as in Michigan) are generally more effective. Figure 7 shows a seats-votes curve for each counterfactual commission scenario (the orange lines), which display the predicted number of total Democratic seats in the US House (y-axis) for a given national Democratic vote share (x-axis). The model predicts that all commission structures would reduce the existing Republican advantage in the House driven by both the inefficient geographic distribution of Democratic voters (Chen and Cottrell,, 2016; Chen and Rodden,, 2015) and net partisan manipulation favoring Republicans (Kenny et al.,, 2023). We also compare these estimated relationships to the actual seats-votes curves for 2010 and 2020 (the black lines). We note that the lines for 2010 and 2020 intersect the 50% vote line below the 50% seat line, since a combination of the underlying geographic distribution of Democratic votes and gerrymandering efforts nationwide disadvantages the Democratic party.
We find that New York- and Ohio-style reforms produce relatively moderate improvements in partisan bias and responsiveness. The point estimates for the number of Democratic seats gained under these reforms is 5.9 for New York-style reforms, with 95% credible interval (–0.39, 12.02), and 8.2 for Ohio-style reforms (1.1, 15.2). Responsiveness for New York- and Ohio-style reforms is also similar as for every 1pp increase in vote share, Democrats gain an estimated 8 seats with New York-style reforms, and 8.1 seats with Ohio-style reforms. These responsiveness values are only slightly higher than the actual responsiveness following 2020 redistricting. Ohio-style reforms appear more effective overall, though the substantive differences are small; we are more than 99% confident that Ohio-style reforms produce a larger effect on partisan bias than New York-style reforms and 86% confident that Ohio-style reforms increase responsiveness more than New York-style reforms do.
Further, we find that Michigan-style reforms have the greatest effects in both responsiveness and Democratic seats, since partisan actors are the most constrained by the presence of a nonpartisan commission, no partisan veto points, and the potential for court review. We estimate that the number of Democratic seats gained under Michigan-style reforms is 22 with 95% credible interval (7.4, 38.3), and that for every 1pp increase in vote share, Democrats gain an estimated 9 seats (6.8, 11.2), an increase of 1.2 over the baseline. We are more than 99% confident that Michigan-style reforms have a larger effect than Ohio reforms, and more than 99% confident in a larger effect than New York-style reforms. We are also 87% confident that Michigan-style reforms increase responsiveness.
Though all three proposed reforms reduce the deviation from partisan symmetry by increasing net Democratic seats, Michigan-style reforms suggest that much greater effects are produced when multiple nodes of the redistricting game constrain leeway by ensuring partisan nodes are nonpartisan, and do not precede partisan vetoes. Meanwhile, examples of New York-style and Ohio-style reforms demonstrate how constraining partisan actors at different nodes, through different reforms, may produce substantively similar effects.
7 Conclusion
In this paper, we propose a methodology for estimating causal effects of complex institutional reforms and apply it to study the impacts of redistricting reforms in the United States. Although redistricting reforms differ across states in important procedural details, we show how to obtain a theoretically informed univariate summary measure of such reforms through a game-theoretic model. We then combine the results of this formal model with a standard causal inference research design strategy to obtain credible causal effect estimates even with a limited sample size. We find that redistricting reforms reduce the partisan bias of enacted plans by constraining the leeway of partisan actors.
Our methodological approach enables us to go beyond the estimation of causal effects. Specifically, we conduct a counterfactual policy analysis to infer the consequences of adopting a series of different redistricting reforms nationwide. We find that adopting redistricting commissions generally reduces the current Republican advantage, but commissions are more effective at reducing bias when they place greater restrictions on partisan actors. For example, a Michigan-style reform (which combines a nonpartisan commission with no partisan veto points) would yield a much greater pro-Democratic effect than commissions adopted in Ohio and New York (which maintain the potential for some partisan control). While we apply this approach to three commission reforms, the same model can be applied to other reforms of interest.
The key idea behind our methodology is also applicable to studies of other complex institutions. For example, most scholars analyze democratic institutions by using the Polity score, which is a 21-point scale index based on six “component variables” concerning executive recruitment, executive constraints, and political participation. Instead of summing these scores to obtain the overall Polity score, our proposed methodology suggests that researchers consider developing a more theoretically informed measure of democracy, based directly on their substantive questions of interest. We believe that such an approach preserves critical aspects of institutional complexity while improving the credibility of causal analysis.
[…]
Below, we provide a simplified summary of the coding, prior to the full details of the coding scheme.
[…]
Drawer.
What kind of body draws the maps?
  • Legislature: The state legislature directly draws the maps. Example: Florida 2020.
  • Commission: Embodies any type of commission, including independent commission, bipartisan commission, or a legislative commission. Example: California 2020 had an independent commission.
  • N/A: There is no redistricting. Example: Wyoming 2020 only has one congressional district and so does not redistrict.
Drawer Control.
Who effectively controls the map-drawing body?
  • Democrats. Example: Nevada 2020 had Democratic majorities in both chambers of the state legislature.
  • Republicans. Example: North Carolina 2020 had Republican majorities in both chambers of the state legislature.
  • Split: The map is drawn by the legislature, and partisan control of the legislature is split across both chambers. Example: Minnesota 2020 had a Democratic majority in the state house and a Republican majority in the state senate, so the legislature was under split partisan control.
  • Nonpartisans: The map-drawing body is composed of people chosen through nonpartisan or explicitly bipartisan means. Example: California 2020 had an independent commission.
  • N/A
Veto 1.
Who has veto power over the proposed maps?
  • Legislature: The legislature can choose to adopt or substitute their own maps for those presented to it by the initial map drawers. Example: Iowa 2020. The legislature can veto the commission’s maps.
  • Governor. Example: Alabama 2020.
  • Voters: Maps can be subject to public referendum. Example: California 2020.
  • N/A: No body has veto authority, or the map drawer has the votes/ability to override vetoes. Example: Kentucky 2020. Though the governor has statutory authority to veto maps from the legislature, the legislature had a Republican supermajority that overrode the Democratic governor’s veto.
Veto 1 Control.
Who controls the body with veto 2 power?
  • Democrats. Example: Illinois 2020. The Democratic governor can veto the legislature’s plan.
  • Republicans. Example: Georgia 2020. The Republican governor can veto the legislature’s plan.
  • Split: The veto power is held by the legislature, which is split in its partisan control across both of its chambers. Example: Washington 2010. The legislature has veto power over the commission’s maps, but the legislature was under split partisan control.
  • N/A: No body has veto authority, or the map drawer has the votes/ability to override vetoes. Example: Tennessee 2020. The legislature has a Republican supermajority, and though the Republican governor has the statutory authority to veto maps, a supermajority renders this power effectively moot.
Veto 2.
Who else has veto power over the proposed maps?
  • Governor. Example: Iowa 2020. After the legislature receives maps from the legislature and approves a set of maps, the governor can veto the redistricting bill.
  • N/A
Veto 2 Control.
Who controls the body with veto 2 power?
  • Democrats. Example: Maine 2020. The Democratic governor can veto the plan from the commission which was approved by the legislator.
  • Republicans. Example: Iowa 2020. The Republican governor can veto the plan from the commission which was approved by the legislator.
  • N/A: No body has veto 2 authority, or the map drawer has the votes/ability to override vetoes. Example: Utah 2020. The legislature (which has veto power over the commission) has a Republican supermajority, and though the Republican governor has the statutory authority to veto maps, a supermajority renders this power effectively moot.
Court Review.
Is there a legal avenue for map challenges to partisan gerrymandering in state court? (Rudensky,, 2023; Wang et al.,, 2019; Douglas,, 2014)
  • Yes: Clear enabling provisions in state law or constitution OR a legal challenge brought at any point in the past which was not dismissed. Example: New Mexico 2020. A state court accepted (but ultimately denied) a lawsuit from the New Mexico GOP challenging Congressional lines as a partisan gerrymander.
  • Maybe: Possible enabling provisions in state law or constitution AND no legal challenge brought at any point. Example: South Carolina 2020.
  • No: No enabling provisions in state law or constitution; or legal challenge clearly dismissed. Example: Kansas 2020. The State Supreme Court ruled that the use of partisan factors in redistricting is constitutionally permissible.
  • N/A
Note that court review is a hidden variable that is not clear until a legal challenge is actually brought forth and either succeeds or is rejected. Indeed, even the specter of litigation may influence how other actors in redistricting approach the drawing of maps. When legal challenges were brought forth for the first time in a state in the 2020 redistricting cycle and the relevant laws under which the court ruled were in place in the 2010 redistricting cycle, we retroactively apply the permissibility of court review to the 2010 cycle as well under the assumption that had a challenge been brought in 2010, it would have proceeded the same way.
Court Control.
Which party controls the state supreme court (or equivalent)? (Parker,, 2023)
  • Democrats. Example: Illinois 2020.
  • Republicans. Example: Indiana 2020.
  • Split: Equally divided court during the window for which challenges could have been brought for 2012 and 2022 maps. Example: Nevada 2020.
  • Nonpartisans: Supreme court is truly nonpartisan, not just having nonpartisan elections. Example: South Carolina 2020.
  • N/A
Stalemate 1.
Who decides the outcome if the initial map drawer has a stalemate?
  • Court: Court has statutory power to intervene in the event of a stalemate. Example: New Jersey 2020.
  • Commission. Example: Oregon 2020.
  • Commission Staff. Example: Colorado 2020.
  • Unclear: no clear statutory or constitutional provision. Example: Louisiana 2020.
  • N/A
Stalemate 1 Control.
Who controls the initial stalemate outcome?
  • Democrats. Example: Maine 2020. The Maine Supreme Court, which can intervene in stalemates, is controlled by Democrats.
  • Republicans. Example: Indiana 2020. The Indiana Supreme Court, which can intervene in stalemates, is controlled by Republicans.
  • Split: Stalemates are broken by a body that must have bipartisan support. Example: Ohio 2020. Stalemates are broken by a backup commission with the support of at least two members from each party (see exceptions below.)
  • Nonpartisans: includes any type of commission selected in a nonpartisan or explicitly bipartisan way. Example: Oregon 2020.
  • N/A
Stalemate 2.
Who decides the outcome if the first stalemate-breaking body itself has a stalemate?
  • Court: Court has statutory power to intervene in the event of a stalemate. Example: Connecticut 2020.
  • Legislature. Example: Ohio 2020.
  • Unclear. Example: Minnesota 2020.
  • N/A
Stalemate 2 Control.
Who controls the second stalemate outcome?
  • Democrats. Example: Connecticut 2020.
  • Republicans. Example: Ohio 2020.
  • N/A
Final Drawer.
Who drew the map in place for the 2012 or 2022 election cycle?
  • Legislature: Includes re-drawing a remedial map after court challenges. Example: South Carolina 2020.
  • Commission: Includes re-drawing a remedial map after court challenges. Example: Washington 2020.
  • Governor. No example available.
  • Court master: Court ordered a special master to draw the final plan. Example: Pennsylvania 2020.
  • Court D remedy: Court picked a remedial map from a Democratic-affiliated plaintiff or intervenor. Example: Colorado 2020.
  • Court R remedy: Court picked a remedial map from a Republican-affiliated plaintiff or intervenor. No example available.
  • N/A
Preclearance.
Was any part of the state subject to DOJ preclearance in 2010 (prior to Shelby County v. Holder)? (U.S. Department of Justice Civil Rights Division,, 2013)
  • Yes. Example: Georgia 2010.
  • No. Example: Massachusetts 2010.
[…]
Most criteria are straightforward, and the information for coding can easily be found in public data for each state. Furthermore, most states have similar redistricting policies that are easily classified under the categories we defined earlier. However, there are unique exceptions that we will treat below. In particular, we coded the below issues as follows.
Veto overrides.
  • In Kansas 2020, Kentucky 2020, Louisiana 2020, Maryland 2010, Maryland 2020, Massachusetts 2010, Massachusetts 2020, Montana 2020, New York 2020, Rhode Island 2010, and Rhode Island 2020, the legislature had the necessary supermajority to override any vetoes, so we coded the governor’s veto as N/A.
Unique commission compositions.
  • In Idaho in 2010 and 2020, commissions are re-formed upon stalemate. Members are picked by both parties in a way that forms a 3-3 even split from both parties, except that the State Supreme Court picks the members of the commission if the parties don’t decide in time. We code the control of this commission as having a partisan split.
  • In Indiana in 2010 and 2020, stalemates are broken by a five-member backup commission composed of the majority leader from each house, the chair of the redistricting committee from each house, and a state legislator appointed by the governor. Since both houses and the governorship were controlled by Republicans, we coded this commission as being controlled by the Republican party.
  • In Michigan 2020, if the commission does not adopt a final plan, a plan is randomly chosen. We code this as a stalemate 1 breaker by the nonpartisan commission.
  • In Ohio 2020, if the legislature fails to redistrict by bipartisan majorities, the work falls to a commission that must approve a map by a bipartisan vote. The commission is composed of appointees by state executive officers and legislative parties. If the commission cannot do so, it can pass a map by a majority vote without bipartisan support, but the map only lasts two election cycles instead. This is encoded as a draw by a split state legislature with the first stalemate broken by a nonpartisan commission and the second stalemate broken by Republicans, since Republicans control all state executive offices and have a functional majority on the commission.
Courts.
  • In Kansas in 2010, the legislature failed to redistrict, and a federal court stepped in to draw the maps instead. But since this was not done under any statutory power and only done once, we consider this an exception to the normal redistricting process in Kansas rather than a procedurally institutionalized stalemate breaker. Similarly in Wisconsin in 2020, the State Supreme Court approved “least change” maps drawn by the governor after a stalemate, but we coded the stalemate as unclear for the same reasons.
  • On the other hand, in Minnesota in 2010 and 2020 we have coded the court as a stalemate breaker because it has consistently intervened in previous stalemates.
  • In Minnesota in 2020, a court-appointed panel drew the final map. We code this as equivalent to that of a court master.
Relevant history that did not affect encoding.
  • In North Carolina, the 2012 map was ultimately struck down but only after the election. Furthermore, the 2022 map was an interim plan only used for the 2022 election.
  • In Virginia in 2010, the state legislature initially deadlocked, but drew a map in time after state legislative elections in 2011. Thus, the stalemate procedure for 2010 was coded as unclear.
[…]
The game tree is specified in Figure 2 and is described in Section 3.2. Here we detail the parametric specification of the moves by nature, starting with the post-enactment court review process. The parameters in the model all receive probabilistic priors, rather than being fixed to specific values. These priors place most of their mass across a broad range of plausible values, based on our experience and observation of past redistricting processes.
[…]
We decompose the court challenge process into five components, as follows:
  1. 1.
    The probability a legal challenge is possible, which depends on the data coding, is specified as
[…]
    The prior on chal_poss_conf is $\mathrm{Beta}\quantity(19,1)$ and the prior on chal_poss_maybe is $\mathrm{Beta}\quantity(6,14)$. The effect of this setup is to follow the treatment coding closely but allow for some uncertainty in the possibility of court review, rather than, e.g., precluding any possibility of court challenge in a state because it was coded as “No.” State courts may change their interpretations of state laws and constitutions in any given case.
  2. 2.
    The probability a challenge is made when one is possible, which depends on the extremity of the partisan bias of the first enacted plan, $x$. Let $F(x)=\pi^{-1}\arctan(x)+\frac{1}{2}$ be the CDF of a standard Cauchy distribution. The specification of this probability is then
[…]
    $a$ $=F^{-1}(\texttt{chal\_prob\_bias}_{0})$
    $b$ $=(F^{-1}(\texttt{chal\_prob\_bias}_{2})-F^{-1}(\texttt{chal\_prob\_bias}_{0}))/2^{2}.$
[…]
  3. 3.
    The probability a court sides with plaintiffs challenging the first enacted plan, i.e., that the court intervenes, which depends on the partisan bias of the first enacted plan as well as the partisan makeup of the state supreme court. We assume that courts with a partisan majority are more likely to intervene and find in favor of challengers when the partisan bias of the first enacted plan is opposite in sign to the court’s partisan lean. Let $g(x;a)=x^{2}(0.7+ac\cdot x+(0.2x)^{2})$, with $c=\sqrt{4(2\cdot 0.7)(12\cdot 0.2^{2})}/6$, which is a convex quartic with asymmetry controlled by a parameter $a$; the value $c$ is set so that the discriminant of the second derivative is positive, i.e., that the quartic is indeed convex. Then the intervention probability is
[…]
    with
    $a$ $=F^{-1}(\texttt{interv\_prob\_bias}_{0})$
    $b$ $=(F^{-1}(\texttt{interv\_prob\_bias}_{2})-F^{-1}(\texttt{interv\_prob\_bias}_{0}))/2^{2}$
[…]
  4. 4.
    The expected remedy a court orders when siding with plaintiffs, which depends on the partisan bias of the first enacted plan as well as the partisan makeup of the state supreme court, and is specified as
[…]
  5. 5.
    The probability and expected effect of a challenge based on the federal Voting Rights Act, which depends on whether a state was subject to DOJ preclearance pre-Shelby, as well as the partisan bias of the first enacted plan. The probability of a VRA challenge is zero for non-preclearance states; for preclearance states it is
[…]
    $a$ $=F^{-1}(\texttt{vra\_chal\_prob\_bias}_{0})$
    $b$ $=(F^{-1}(\texttt{vra\_chal\_prob\_bias}_{2})-F^{-1}(\texttt{chal\_prob\_bias}_{0}))/2.$
[…]
    We assume the probability of a successful VRA claim does not depend on the partisan bias of the first enacted plan; this parameter is vra_interv_prob and receives a prior of $\mathrm{Beta}\quantity(4,1.5)$. The court-ordered remedy for a successful VRA challenge is specified as
[…]
Putting these five pieces together, we can write the expected outcome of the post-enactment processes as
[…]
When there is no stalemate process specified, when stalemate processes are exhausted, or when the stalemate process is controlled by neither party, the stalemate outcome is assumed to be a linear rescaling of the bias of the plan drawn at the previous stage, $x$, which acts as a sort of anchoring point, and the partisan control of the initial map drawer. The specification is
[…]
with out_nonp_part_adv defined as above, and stale_slope receiving a $\mathrm{Beta}\quantity(3,17)$ prior. This reflects a status quo bias towards the last-drawn plan.
When stalemates are specified as being resolved by state courts, and the state court is partisan, the stalemate outcome is
$\begin{cases*}\texttt{stale\_slope}\cdot x-\texttt{out\_nonp\_part\_adv}&if Court Control is coded as ``Democrats''\\ \texttt{stale\_slope}\cdot x+\texttt{out\_nonp\_part\_adv}&if Court Control iscoded as ``Republicans''\end{cases*}$
given an input partisan bias $x$.
[…]
When veto control is coded as NA (other than in the case of a governor with overridable veto) or a court has a veto, the veto is exercised with probability
[…]
with the parameter veto_nonp_prob_max receiving a $\mathrm{Beta}\quantity(3,7)$ prior. So these veto players are unlikely to intervene, but relatively more likely if the proposed plan is more extreme
[…]
Map drawers with split partisan control are assumed to stalemate with probability controlled by parameter stale_split_prob, which has a $\mathrm{Beta}\quantity(3,5)$ prior. When they don’t stalemate, split-control map drawers are assumed to propose a map with the same partisan bias as one drawn by a nonpartisan map drawer. While we suspect this probability could be set to 1 without changing the results, a probability of compromise reflects a history of compromise plans in state legislatures.
In Round 1, nonpartisan map drawers are assumed to propose a plan with partisan bias of 0. In Round 2—after one veto has already occurred—nonpartisan map drawers are assumed to shift the bias of their proposed plan in the direction of the average partisanship of the veto players. Specifically, if $x$ is the proposal from the previous stage, then the proposal in Round 2 is
[…]
The veto_nonp_shift parameter controls the amount of the shift and receives a $\mathcal{N}\quantity(0.65,0.3)$ prior. This mechanism reflects real-world patterns of repeated map-drawing, such as when the (Republican-controlled) Iowa legislature rejected the proposed plan from the nonpartisan advisory commission, then adopted a second proposal from the same commission that favored Republicans. more.
[…]
In Alabama, the legislature draws congressional districts, subject to a governor’s veto, and there is no mechanism for state court review. Thus, Alabama’s process would be described by the “Drawer” and “Veto 1” steps only in Round 1 of the Figure 2 process.
The first move belongs to the Republican party, which controls the state legislature. It has to pick the amount of partisan bias $-4<x<4$ in the plan that it adopts. If this plan is ultimately the final enacted plan, the Republicans receive utility $x$ and the Democrats receive utility $-x$. The second move also belongs to the Republican party, which controls the governorship. The party must decide whether or not to veto the adopted plan; if it does, there is a second round of drawing and vetoing by the legislature and governor (controlled by Republicans).
If a plan is adopted at either the first or second round, it proceeds to possible court review. The court may decide to accept a legal challenge, decide in favor of the plaintiffs, and redraw the map; this choice is considered a move by nature. If the court review results in a redrawn map with partisan bias $x^{\prime}$, then the Republicans receive utility $x^{\prime}$ and the Democrats receive utility $-x^{\prime}$. In Alabama, court review is generally not allowed on partisan grounds, but challenges under the federal Voting Rights Act are possible, so there is moderate probability of the legislature-adopted plan being overturned. If the legislature’s plan is vetoed both times, the process results in a stalemate. Since there is no enumerated stalemate procedure in Alabama, courts must step in and redraw district lines to ensure compliance with federal constitutional “one person, one vote” apportionment requirements. This is also considered a move by nature.
To calculate the equilibrium itself, we numerically solve the game via backwards induction. This requires up to four levels of nested optimization. For Alabama, this means that we start with the last partisan move in the game, which is the Round 2, Veto 1 player, the Republican governor. She must decide, given a proposed partisan bias $x$ from the Round 2 map drawer, whether to veto the plan, in which case the enacted plan is determined via the court stalemate process, or to accept the plan and let it become law. In either case, the resulting plan may be challenged in court.
[…]
For completeness, we consider a series of alternative outcomes which measure the partisan bias of redistricting plans. We first consider an alternative to the efficiency gap, the dilution asymmetry (Gordon and Yntiso,, 2024). This measures the difference in the percentage of each party’s votes that are wasted (cf. the number of wasted votes used in the calculation of the efficiency gap). Similar to the efficiency gap, we consider both the absolute value of differences from simulations and signed raw values for the dilution asymmetry.
A series of partisan bias measures are based on the seats-votes curve. Under a fair plan, the seats-votes curve should be symmetric (Katz et al.,, 2020). We consider two groups of measures based on the seats-votes curve. First, we consider the “partisan bias” which measures the deviation from symmetry in the seats-votes curve. We use both the raw signed value and absolute value of the difference from simulations of the partisan bias.
Finally, we consider total and differential individual harm, which measures how a redistricting plan negatively affects the ability for a Democratic or Republican voter to elect a candidate of their choice, relative to simulated redistricting plans (McCartan and Kenny,, 2022). Differential partisan harm is a signed measure where positive scores indicate Democrats are harmed at a higher rate than Republican voters. The total partisan harm of a plan is a nonpartisan measure, which represents the total harm to both parties.
[…]
To evaluate robustness to violations of the sufficiency assumption, i.e., that
$Y_{it}(\mathbf{z})=Y_{it}(\mathbf{z}^{\prime})\mbox{\quad for any\quad}\mathbf{z},\mathbf{z}^{\prime}\text{ with }u^{*}(\mathbf{z})=u^{*}(\mathbf{z}^{\prime}),$
we re-run the ACR estimates including $\mathbf{z}$ as a covariate. If the assumption were to hold, $\mathbf{z}$ should have no effect on the outcome, since $u$ is controlled for. Of course, the data are limited ($n=43$), and the dimension of $\mathbf{z}$ is relatively large—the six elements we use here must be included as levels and as differences, and are each interacted with the change in $u$. Thus we would likely expect some changes due solely to random noise.
Some of the elements of $\mathbf{z}$ do not change from 2010-2020 or exhibit variation in only one or two states. We remove these elements, and re-code several “control” variables as Democratic/Republican/Other to further avoid exact collinearity. The final specification for additional covariates included in ACR estimation is as follows:
  • drawer in 2010
  • Indicator for drawer_ctrl being democrats in 2010
  • Indicator for drawer_ctrl being republicans in 2010
  • Indicator for veto_1 being governor in 2010
  • Indicator for veto_1_ctrl being democrats in 2010
  • Indicator for veto_1_ctrl being republicans in 2010
  • court_review in 2010
  • Indicator for court_ctrl being democrats in 2010
  • Indicator for court_ctrl being republicans in 2010
  • Indicator for any change in drawer from 2010 to 2020
  • Indicator for any change in drawer_ctrl from 2010 to 2020
  • Indicator for any change in veto_1 from 2010 to 2020
  • Indicator for any change in veto_1_ctrl from 2010 to 2020
  • Indicator for any change in court_ctrl from 2010 to 2020
[…]
Table LABEL:tbl-coefs below contains point estimates and standard errors for all estimated model coefficients. Each set of rows marked by a single value in the “Model” column represents a separately estimated model. “D.2” in the “Spec.” (specification) column indicates the robustness check specification used in the corresponding appendix.
[…]
Model Spec. Leeway Outcome type Outcome Coefficient Estimate Std. err.
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst 0.09221 0.16980
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. nash_worst_10 0.12040 0.03761
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dvs_08 -0.05920 0.30000
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. is_south 0.19200 0.14240
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. log(n_distr_20) 0.01992 0.13900
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. distr_chg 0.57970 0.13120
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. log(1 + corrupt) -0.15150 0.10380
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. initiative 0.14910 0.10740
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. drawer_ctrl_chg -0.32900 0.14020
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. court_ctrl_chg -0.18110 0.10700
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:nash_worst_10 -0.00306 0.02215
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:dvs_08 -0.00818 0.15090
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:is_south 0.00017 0.10790
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:log(n_distr_20) 0.00404 0.03384
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:distr_chg 0.02896 0.08761
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:log(1 + corrupt) 0.00671 0.02437
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:initiative 0.04111 0.08927
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:drawer_ctrl_chg 0.01838 0.07799
1a Main Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:court_ctrl_chg -0.05239 0.08557
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst 0.11580 0.24050
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. nash_worst_10 0.17760 0.09216
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dvs_08 -0.07825 0.83470
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. is_south 0.26910 0.17460
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. log(n_distr_20) -0.01311 0.19660
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. distr_chg 0.61650 0.17400
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. log(1 + corrupt) -0.08925 0.12210
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. initiative 0.11810 0.15210
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. drawer_10commission 0.02661 0.42290
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. drawer_10legislature -0.12930 0.50660
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE -0.02996 0.26610
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE -0.07491 0.23890
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_10 == "governor")TRUE 0.06005 0.33320
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.29160 0.18320
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.12340 0.20850
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. court_review_10no -0.13350 0.19640
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. court_review_10yes -0.04109 0.24110
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. I(court_ctrl_10 == "democrats")TRUE -0.02576 0.20330
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.06823 0.19430
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. drawer_chg 0.66920 0.61180
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. drawer_ctrl_chg -0.61060 0.22820
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. veto_1_chg -0.17860 0.55080
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. veto_1_ctrl_chg -0.00788 0.19490
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. court_ctrl_chg -0.16390 0.15200
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:nash_worst_10 -0.00189 0.02305
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:dvs_08 0.00668 0.14860
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:is_south -0.01019 0.12990
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:log(n_distr_20) 0.00586 0.03286
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:distr_chg 0.01317 0.09716
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:log(1 + corrupt) 0.00589 0.02460
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:initiative 0.01612 0.10420
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:drawer_10legislature 0.00918 0.08789
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:I(drawer_ctrl_10 == "democrats")TRUE -0.11220 0.24190
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:I(drawer_ctrl_10 == "republicans")TRUE 0.00333 0.09171
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:I(veto_1_10 == "governor")TRUE 0.01256 0.08874
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:I(veto_1_ctrl_10 == "democrats")TRUE 0.07221 0.17270
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:I(veto_1_ctrl_10 == "republicans")TRUE -0.00486 0.08932
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:court_review_10no -0.01378 0.09600
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:court_review_10yes 0.10770 0.25920
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:I(court_ctrl_10 == "democrats")TRUE -0.00349 0.19560
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:I(court_ctrl_10 == "republicans")TRUE 0.00849 0.10050
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:drawer_chg -0.00170 0.08706
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:drawer_ctrl_chg 0.03420 0.08638
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:veto_1_chg 0.03763 0.11460
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:veto_1_ctrl_chg -0.02037 0.11280
1a+ D.2 Maximum Nonpartisan Abs. Rep. seat diff. from sims. dose_worst:court_ctrl_chg -0.03574 0.10380
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party -0.02953 0.07412
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. nash_party_10 0.00884 0.03512
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dvs_08 0.31510 0.35620
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. is_south 0.24230 0.16600
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. log(n_distr_20) -0.04537 0.18160
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. distr_chg 0.52630 0.16920
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. log(1 + corrupt) -0.09801 0.11630
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. initiative -0.00210 0.12210
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:nash_party_10 0.01477 0.00977
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:dvs_08 0.00077 0.07767
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:is_south -0.00902 0.04154
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:log(n_distr_20) 0.01839 0.02278
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:distr_chg 0.00832 0.04191
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:log(1 + corrupt) 0.00776 0.01426
1b Main Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:initiative 0.02356 0.04765
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party -0.01849 0.10580
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. nash_party_10 0.05474 0.06842
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dvs_08 -0.52070 0.83270
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. is_south 0.24360 0.18610
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. log(n_distr_20) 0.03578 0.18120
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. distr_chg 0.64220 0.16960
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. log(1 + corrupt) -0.12630 0.12310
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. initiative -0.02825 0.16690
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_10commission 0.33330 0.45620
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_10legislature 0.28390 0.45460
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.50630 0.32010
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.11920 0.28790
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_10 == "governor")TRUE 0.04854 0.31250
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.26510 0.21340
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.23410 0.21230
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. court_review_10no 0.05293 0.17120
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. court_review_10yes -0.16650 0.20670
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.08932 0.20230
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.06643 0.20780
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_chg 0.42410 0.48240
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_ctrl_chg -0.40460 0.25670
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. veto_1_chg -0.34770 0.49490
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. veto_1_ctrl_chg 0.16270 0.20950
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. court_ctrl_chg -0.19570 0.14800
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:nash_party_10 0.00692 0.01309
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:dvs_08 -0.00697 0.08112
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:is_south 0.01534 0.04504
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:log(n_distr_20) 0.00972 0.02290
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:distr_chg 0.00286 0.05057
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:log(1 + corrupt) 0.00500 0.01448
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:initiative 0.03453 0.05159
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:drawer_10legislature 0.00068 0.03873
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:I(drawer_ctrl_10 == "democrats")TRUE -0.01115 0.04667
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:I(drawer_ctrl_10 == "republicans")TRUE 0.00334 0.06387
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:I(veto_1_10 == "governor")TRUE -0.00096 0.04059
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:I(veto_1_ctrl_10 == "democrats")TRUE -0.03222 0.05892
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:I(veto_1_ctrl_10 == "republicans")TRUE 0.06267 0.06431
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:court_review_10no -0.01466 0.04219
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:court_review_10yes -0.08544 0.22360
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:I(court_ctrl_10 == "democrats")TRUE -0.01829 0.04838
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:I(court_ctrl_10 == "republicans")TRUE 0.04300 0.07919
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:drawer_chg 0.00769 0.09602
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:drawer_ctrl_chg 0.00920 0.04067
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:veto_1_chg 0.04228 0.11570
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:veto_1_ctrl_chg -0.01118 0.05015
1b+ D.2 Realized Nonpartisan Abs. Rep. seat diff. from sims. dose_party:court_ctrl_chg -0.02578 0.04226
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty 0.02352 0.10360
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. nash_absparty_10 0.10540 0.03964
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dvs_08 -0.06296 0.29670
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. is_south 0.08922 0.15330
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. log(n_distr_20) 0.04741 0.14340
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. distr_chg 0.48780 0.12850
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. log(1 + corrupt) -0.14970 0.10630
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. initiative -0.02183 0.11800
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:nash_absparty_10 0.01907 0.02148
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:dvs_08 0.00208 0.11190
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:is_south 0.00757 0.06489
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:log(n_distr_20) 0.00666 0.02756
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:distr_chg 0.02649 0.07483
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:log(1 + corrupt) 0.00818 0.01918
1c Main Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:initiative 0.05726 0.06295
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty 0.10510 0.14740
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. nash_absparty_10 0.12150 0.06606
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dvs_08 -0.10220 0.77560
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. is_south 0.14700 0.18000
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. log(n_distr_20) 0.02520 0.17810
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. distr_chg 0.55750 0.14820
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. log(1 + corrupt) -0.08689 0.12200
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. initiative -0.02780 0.16350
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_10commission -0.05704 0.43340
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_10legislature 0.04620 0.44580
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.18770 0.35310
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.28670 0.30800
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_10 == "governor")TRUE -0.25630 0.34810
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.36110 0.18390
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.22370 0.18630
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. court_review_10no -0.00425 0.15750
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. court_review_10yes -0.07536 0.20260
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.11440 0.19240
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. I(court_ctrl_10 == "republicans")TRUE 0.02976 0.17620
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_chg 0.64930 0.49310
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. drawer_ctrl_chg -0.55580 0.21720
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. veto_1_chg -0.34080 0.50540
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. veto_1_ctrl_chg 0.22110 0.18320
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. court_ctrl_chg -0.25520 0.14120
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:nash_absparty_10 0.00661 0.02399
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:dvs_08 0.00309 0.10370
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:is_south 0.07035 0.07945
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:log(n_distr_20) 0.00815 0.02711
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:distr_chg 0.01300 0.08689
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:log(1 + corrupt) 0.00960 0.01900
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:initiative 0.05460 0.06985
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:drawer_10legislature 0.00021 0.05664
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:I(drawer_ctrl_10 == "democrats")TRUE 0.01149 0.12750
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:I(drawer_ctrl_10 == "republicans")TRUE -0.01515 0.06927
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:I(veto_1_10 == "governor")TRUE -0.00024 0.06012
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:I(veto_1_ctrl_10 == "democrats")TRUE -0.11190 0.09327
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:I(veto_1_ctrl_10 == "republicans")TRUE 0.04110 0.06466
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:court_review_10no -0.01955 0.06058
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:court_review_10yes 0.20820 0.21170
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:I(court_ctrl_10 == "democrats")TRUE -0.03767 0.09785
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:I(court_ctrl_10 == "republicans")TRUE 0.02482 0.07356
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:drawer_chg -0.03751 0.09705
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:drawer_ctrl_chg 0.01761 0.07356
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:veto_1_chg 0.00421 0.11910
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:veto_1_ctrl_chg -0.02425 0.06912
1c+ D.2 Abs. realized Nonpartisan Abs. Rep. seat diff. from sims. dose_absparty:court_ctrl_chg -0.03878 0.06647
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst -0.17330 0.94980
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. nash_worst_10 0.74240 0.24330
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dvs_08 -0.51310 1.94600
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. is_south 0.11750 0.93770
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. log(n_distr_20) -0.27490 0.90250
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. distr_chg 2.08600 0.90520
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. log(1 + corrupt) -0.51050 0.66760
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. initiative 0.87420 0.71780
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. drawer_ctrl_chg -1.70800 0.91610
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. court_ctrl_chg -1.05100 0.70810
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:nash_worst_10 0.00258 0.12930
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:dvs_08 -0.13900 0.84040
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:is_south 0.43490 0.61120
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:log(n_distr_20) 0.05172 0.19080
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:distr_chg 0.08939 0.50310
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:log(1 + corrupt) 0.05179 0.13770
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:initiative 0.01218 0.53200
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:drawer_ctrl_chg 0.34890 0.46750
2a Main Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:court_ctrl_chg -0.08691 0.49540
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst -0.02887 1.33700
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. nash_worst_10 1.22300 0.53070
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dvs_08 -1.33200 4.68400
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. is_south 0.32050 0.99100
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. log(n_distr_20) 0.15000 1.05400
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. distr_chg 2.25700 0.91860
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. log(1 + corrupt) -0.38770 0.73110
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. initiative 0.62000 0.91120
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. drawer_10commission 0.89900 2.42200
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. drawer_10legislature -1.06600 2.79300
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. I(drawer_ctrl_10 == "democrats")TRUE -0.11260 1.49900
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. I(drawer_ctrl_10 == "republicans")TRUE -0.82140 1.40200
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_10 == "governor")TRUE 0.41770 1.96700
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_ctrl_10 == "democrats")TRUE 1.76400 1.09000
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.36110 1.22900
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. court_review_10no -1.01500 1.07900
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. court_review_10yes -1.16400 1.31500
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. I(court_ctrl_10 == "democrats")TRUE -0.51480 1.10500
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. I(court_ctrl_10 == "republicans")TRUE -1.59000 1.12600
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. drawer_chg 1.15400 3.77000
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. drawer_ctrl_chg -2.45500 1.27400
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. veto_1_chg -0.15250 3.47500
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. veto_1_ctrl_chg -0.71440 1.10500
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. court_ctrl_chg -0.71840 0.89330
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:nash_worst_10 -0.00232 0.13580
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:dvs_08 -0.01747 0.89250
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:is_south 0.17290 0.68670
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:log(n_distr_20) 0.01699 0.20010
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:distr_chg 0.01910 0.52540
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:log(1 + corrupt) 0.01884 0.14210
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:initiative -0.05804 0.61950
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:drawer_10legislature -0.01654 0.49450
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:I(drawer_ctrl_10 == "democrats")TRUE -1.57300 1.38500
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:I(drawer_ctrl_10 == "republicans")TRUE 0.09715 0.46180
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:I(veto_1_10 == "governor")TRUE 0.05743 0.46600
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:I(veto_1_ctrl_10 == "democrats")TRUE 0.13780 0.99840
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:I(veto_1_ctrl_10 == "republicans")TRUE 0.02647 0.50490
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:court_review_10no 0.06764 0.49430
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:court_review_10yes 0.44170 1.43900
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:I(court_ctrl_10 == "democrats")TRUE -0.56720 1.13400
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:I(court_ctrl_10 == "republicans")TRUE -0.00322 0.56740
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:drawer_chg 0.11950 0.52100
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:drawer_ctrl_chg 0.22460 0.47940
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:veto_1_chg 0.23090 0.64610
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:veto_1_ctrl_chg -0.07623 0.66160
2a+ D.2 Maximum Nonpartisan Abs. Rep. seats z-score from sims. dose_worst:court_ctrl_chg -0.16290 0.56580
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party -0.11060 0.43890
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. nash_party_10 -0.03099 0.20950
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dvs_08 1.85100 2.19900
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. is_south 0.48280 1.06900
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. log(n_distr_20) -0.52510 1.06700
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. distr_chg 1.85700 0.99820
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. log(1 + corrupt) -0.25160 0.68840
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. initiative 0.11560 0.76700
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:nash_party_10 0.07642 0.05803
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:dvs_08 -0.00335 0.44340
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:is_south 0.01616 0.24410
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:log(n_distr_20) 0.07792 0.13320
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:distr_chg -0.02768 0.25750
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:log(1 + corrupt) 0.03807 0.08180
2b Main Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:initiative 0.04658 0.26430
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party -0.09383 0.58190
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. nash_party_10 0.38280 0.41000
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dvs_08 -1.26600 4.87900
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. is_south 0.47860 1.10900
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. log(n_distr_20) 0.08232 1.07700
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. distr_chg 2.58000 1.02300
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. log(1 + corrupt) -0.44780 0.73170
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. initiative -0.11100 1.07400
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_10commission 1.98100 2.65600
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_10legislature 1.55700 2.66900
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. I(drawer_ctrl_10 == "democrats")TRUE 2.97200 2.00800
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. I(drawer_ctrl_10 == "republicans")TRUE -0.48000 1.75900
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_10 == "governor")TRUE -0.21150 1.85500
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_ctrl_10 == "democrats")TRUE 1.70500 1.29100
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.07465 1.22300
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. court_review_10no -0.76370 1.06300
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. court_review_10yes -2.80800 1.29400
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. I(court_ctrl_10 == "democrats")TRUE 0.49760 1.18200
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. I(court_ctrl_10 == "republicans")TRUE -1.18900 1.18500
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_chg 1.97900 2.92400
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_ctrl_chg -1.52700 1.50000
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. veto_1_chg -1.10600 3.04100
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. veto_1_ctrl_chg 0.93320 1.29300
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. court_ctrl_chg -0.70180 0.90630
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:nash_party_10 0.03681 0.07148
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:dvs_08 -0.06328 0.44690
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:is_south 0.08851 0.25500
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:log(n_distr_20) 0.03982 0.13560
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:distr_chg -0.01794 0.27750
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:log(1 + corrupt) 0.02469 0.08550
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:initiative 0.05302 0.28270
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:drawer_10legislature 0.00959 0.21280
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:I(drawer_ctrl_10 == "democrats")TRUE -0.04754 0.27620
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:I(drawer_ctrl_10 == "republicans")TRUE 0.11610 0.37270
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:I(veto_1_10 == "governor")TRUE 0.01412 0.22640
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:I(veto_1_ctrl_10 == "democrats")TRUE -0.18790 0.34830
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:I(veto_1_ctrl_10 == "republicans")TRUE 0.35750 0.35800
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:court_review_10no -0.00002 0.24650
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:court_review_10yes -0.62900 1.27000
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:I(court_ctrl_10 == "democrats")TRUE -0.15560 0.27700
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:I(court_ctrl_10 == "republicans")TRUE 0.13930 0.43520
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:drawer_chg 0.12780 0.57550
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:drawer_ctrl_chg 0.05568 0.21410
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:veto_1_chg 0.44540 0.70850
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:veto_1_ctrl_chg 0.00678 0.29450
2b+ D.2 Realized Nonpartisan Abs. Rep. seats z-score from sims. dose_party:court_ctrl_chg -0.10460 0.23310
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty 0.37500 0.56850
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. nash_absparty_10 0.64550 0.25560
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dvs_08 -0.28410 1.92500
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. is_south -0.58200 0.91960
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. log(n_distr_20) -0.05096 0.85670
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. distr_chg 1.82900 0.86120
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. log(1 + corrupt) -0.50680 0.64040
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. initiative -0.04093 0.68570
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:nash_absparty_10 0.09614 0.11760
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:dvs_08 0.03403 0.57960
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:is_south 0.01762 0.39650
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:log(n_distr_20) 0.04503 0.15010
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:distr_chg -0.06996 0.41900
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:log(1 + corrupt) 0.04247 0.11180
2c Main Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:initiative 0.24940 0.34320
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty 0.59150 0.85080
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. nash_absparty_10 0.62520 0.42210
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dvs_08 0.25600 4.85500
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. is_south -0.29750 1.15100
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. log(n_distr_20) 0.07921 1.09200
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. distr_chg 2.22100 0.91970
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. log(1 + corrupt) -0.23690 0.73460
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. initiative -0.25350 1.05700
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_10commission 0.47800 2.60900
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_10legislature 0.71780 2.63100
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.77620 2.07500
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.47580 1.79700
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_10 == "governor")TRUE -1.44400 2.14000
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_ctrl_10 == "democrats")TRUE 1.85400 1.19200
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. I(veto_1_ctrl_10 == "republicans")TRUE 0.08277 1.16300
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. court_review_10no -1.00900 1.00100
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. court_review_10yes -2.13300 1.29500
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. I(court_ctrl_10 == "democrats")TRUE 0.79040 1.18400
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. I(court_ctrl_10 == "republicans")TRUE -0.88770 1.15100
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_chg 2.76900 2.89000
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. drawer_ctrl_chg -2.14200 1.32200
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. veto_1_chg -0.83460 3.06000
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. veto_1_ctrl_chg 1.10600 1.15900
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. court_ctrl_chg -1.19700 0.85880
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:nash_absparty_10 0.04077 0.13990
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:dvs_08 0.05560 0.61680
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:is_south 0.26560 0.45890
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:log(n_distr_20) 0.02527 0.15100
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:distr_chg -0.05277 0.50210
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:log(1 + corrupt) 0.03427 0.10900
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:initiative 0.13720 0.40960
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:drawer_10legislature 0.02304 0.32570
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:I(drawer_ctrl_10 == "democrats")TRUE 0.14660 0.76410
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:I(drawer_ctrl_10 == "republicans")TRUE -0.07399 0.40940
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:I(veto_1_10 == "governor")TRUE 0.04241 0.31870
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:I(veto_1_ctrl_10 == "democrats")TRUE -0.58040 0.53070
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:I(veto_1_ctrl_10 == "republicans")TRUE 0.22790 0.38340
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:court_review_10no -0.05994 0.33030
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:court_review_10yes 1.08400 1.22400
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:I(court_ctrl_10 == "democrats")TRUE -0.08218 0.57650
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:I(court_ctrl_10 == "republicans")TRUE 0.08067 0.45810
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:drawer_chg -0.05459 0.54090
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:drawer_ctrl_chg 0.09591 0.41520
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:veto_1_chg 0.26400 0.71980
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:veto_1_ctrl_chg 0.04820 0.40240
2c+ D.2 Abs. realized Nonpartisan Abs. Rep. seats z-score from sims. dose_absparty:court_ctrl_chg -0.28100 0.37720
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst 0.00341 0.01541
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. nash_worst_10 0.01096 0.00350
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dvs_08 -0.00821 0.02781
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. is_south 0.01219 0.01341
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. log(n_distr_20) 0.01801 0.01211
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. distr_chg 0.02947 0.01249
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. log(1 + corrupt) -0.02479 0.00907
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. initiative 0.02275 0.01014
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. drawer_ctrl_chg -0.01387 0.01223
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. court_ctrl_chg -0.01517 0.00978
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:nash_worst_10 -0.00038 0.00183
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:dvs_08 -0.00107 0.01260
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:is_south 0.00058 0.00992
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:log(n_distr_20) 0.00057 0.00290
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:distr_chg 0.00247 0.00798
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:log(1 + corrupt) 0.00058 0.00216
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:initiative 0.00421 0.00816
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:drawer_ctrl_chg 0.00588 0.00708
3a Main Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:court_ctrl_chg -0.00637 0.00782
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst 0.00693 0.02123
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. nash_worst_10 0.02004 0.00740
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dvs_08 -0.05017 0.07315
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. is_south 0.01887 0.01458
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. log(n_distr_20) 0.01893 0.01699
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. distr_chg 0.03983 0.01460
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. log(1 + corrupt) -0.02303 0.01067
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. initiative 0.01762 0.01254
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. drawer_10commission 0.03740 0.03729
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. drawer_10legislature -0.00196 0.04443
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE -0.00622 0.02280
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE -0.01358 0.02112
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_10 == "governor")TRUE 0.00465 0.02700
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.01484 0.01599
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.01013 0.01775
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. court_review_10no -0.01523 0.01645
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. court_review_10yes -0.01201 0.02005
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. I(court_ctrl_10 == "democrats")TRUE -0.00240 0.01635
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.01340 0.01667
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. drawer_chg 0.05746 0.05517
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. drawer_ctrl_chg -0.03249 0.01818
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. veto_1_chg 0.01934 0.05068
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. veto_1_ctrl_chg -0.00317 0.01719
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. court_ctrl_chg -0.00915 0.01341
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:nash_worst_10 -0.00006 0.00202
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:dvs_08 0.00145 0.01319
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:is_south 0.00058 0.01097
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:log(n_distr_20) 0.00046 0.00299
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:distr_chg 0.00101 0.00803
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:log(1 + corrupt) 0.00055 0.00228
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:initiative 0.00165 0.00898
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:drawer_10legislature 0.00006 0.00760
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:I(drawer_ctrl_10 == "democrats")TRUE -0.02031 0.02136
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:I(drawer_ctrl_10 == "republicans")TRUE -0.00035 0.00740
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:I(veto_1_10 == "governor")TRUE 0.00198 0.00727
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:I(veto_1_ctrl_10 == "democrats")TRUE 0.01085 0.01436
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:I(veto_1_ctrl_10 == "republicans")TRUE -0.00075 0.00836
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:court_review_10no 0.00122 0.00779
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:court_review_10yes 0.01879 0.02243
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:I(court_ctrl_10 == "democrats")TRUE -0.00149 0.01622
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:I(court_ctrl_10 == "republicans")TRUE -0.00101 0.00837
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:drawer_chg 0.00246 0.00814
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:drawer_ctrl_chg 0.00557 0.00751
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:veto_1_chg 0.00769 0.01033
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:veto_1_ctrl_chg -0.00406 0.00999
3a+ D.2 Maximum Nonpartisan Abs. efficiency gap, diff. from sims. dose_worst:court_ctrl_chg -0.00500 0.00847
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party 0.00048 0.00680
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. nash_party_10 0.00311 0.00288
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dvs_08 0.02554 0.03165
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. is_south 0.01827 0.01554
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. log(n_distr_20) 0.00976 0.01515
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. distr_chg 0.02779 0.01417
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. log(1 + corrupt) -0.01955 0.01038
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. initiative 0.01592 0.01129
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:nash_party_10 0.00086 0.00089
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:dvs_08 -0.00068 0.00694
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:is_south 0.00016 0.00365
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:log(n_distr_20) 0.00109 0.00205
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:distr_chg 0.00006 0.00385
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:log(1 + corrupt) 0.00049 0.00128
3b Main Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:initiative 0.00014 0.00394
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party -0.00011 0.00872
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. nash_party_10 0.01020 0.00659
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dvs_08 -0.03829 0.07202
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. is_south 0.02641 0.01853
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. log(n_distr_20) 0.01724 0.01638
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. distr_chg 0.03890 0.01550
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. log(1 + corrupt) -0.02143 0.01147
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. initiative 0.01189 0.01626
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_10commission 0.04615 0.04251
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_10legislature 0.03374 0.04274
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.05567 0.03079
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE -0.00104 0.02704
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_10 == "governor")TRUE -0.00887 0.03107
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.01428 0.02091
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.01822 0.01947
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. court_review_10no -0.00308 0.01582
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. court_review_10yes -0.03397 0.01870
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.00217 0.01792
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.01129 0.01880
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_chg 0.03535 0.04414
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_ctrl_chg -0.01648 0.02369
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. veto_1_chg -0.00190 0.04476
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. veto_1_ctrl_chg 0.01267 0.01788
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. court_ctrl_chg -0.00911 0.01309
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:nash_party_10 0.00056 0.00106
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:dvs_08 -0.00033 0.00691
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:is_south 0.00164 0.00379
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:log(n_distr_20) 0.00056 0.00199
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:distr_chg -0.00085 0.00433
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:log(1 + corrupt) 0.00043 0.00121
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:initiative 0.00067 0.00434
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:drawer_10legislature -0.00026 0.00325
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:I(drawer_ctrl_10 == "democrats")TRUE -0.00126 0.00416
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:I(drawer_ctrl_10 == "republicans")TRUE 0.00309 0.00583
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:I(veto_1_10 == "governor")TRUE -0.00031 0.00344
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:I(veto_1_ctrl_10 == "democrats")TRUE -0.00280 0.00540
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:I(veto_1_ctrl_10 == "republicans")TRUE 0.00383 0.00577
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:court_review_10no 0.00044 0.00364
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:court_review_10yes -0.02220 0.02041
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:I(court_ctrl_10 == "democrats")TRUE -0.00173 0.00437
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:I(court_ctrl_10 == "republicans")TRUE 0.00227 0.00671
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:drawer_chg 0.00281 0.00878
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:drawer_ctrl_chg 0.00103 0.00360
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:veto_1_chg 0.00568 0.01109
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:veto_1_ctrl_chg -0.00074 0.00456
3b+ D.2 Realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_party:court_ctrl_chg -0.00151 0.00356
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty 0.00558 0.00885
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. nash_absparty_10 0.01117 0.00368
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dvs_08 -0.00592 0.02871
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. is_south 0.00607 0.01297
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. log(n_distr_20) 0.01719 0.01299
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. distr_chg 0.02705 0.01109
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. log(1 + corrupt) -0.02324 0.00949
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. initiative 0.01142 0.00980
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:nash_absparty_10 0.00133 0.00186
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:dvs_08 0.00132 0.00913
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:is_south -0.00186 0.00568
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:log(n_distr_20) 0.00064 0.00234
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:distr_chg 0.00016 0.00621
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:log(1 + corrupt) 0.00049 0.00158
3c Main Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:initiative 0.00496 0.00486
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty 0.01208 0.01312
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. nash_absparty_10 0.01377 0.00572
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dvs_08 -0.05329 0.06873
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. is_south 0.00877 0.01532
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. log(n_distr_20) 0.02173 0.01545
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. distr_chg 0.02785 0.01291
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. log(1 + corrupt) -0.01971 0.01037
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. initiative 0.00245 0.01359
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_10commission 0.03441 0.04086
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_10legislature 0.02635 0.04079
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.02464 0.03041
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.02117 0.02683
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_10 == "governor")TRUE -0.03230 0.02893
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.01301 0.01550
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.01717 0.01649
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. court_review_10no -0.00652 0.01388
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. court_review_10yes -0.02075 0.01700
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.01183 0.01618
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.00303 0.01591
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_chg 0.06642 0.04209
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. drawer_ctrl_chg -0.02484 0.01874
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. veto_1_chg -0.01048 0.04355
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. veto_1_ctrl_chg 0.01729 0.01604
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. court_ctrl_chg -0.02051 0.01185
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:nash_absparty_10 0.00054 0.00214
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:dvs_08 0.00104 0.00962
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:is_south 0.00452 0.00683
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:log(n_distr_20) 0.00003 0.00253
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:distr_chg -0.00020 0.00732
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:log(1 + corrupt) 0.00020 0.00168
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:initiative 0.00433 0.00609
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:drawer_10legislature 0.00096 0.00486
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:I(drawer_ctrl_10 == "democrats")TRUE 0.00583 0.01132
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:I(drawer_ctrl_10 == "republicans")TRUE -0.00038 0.00605
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:I(veto_1_10 == "governor")TRUE 0.00071 0.00510
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:I(veto_1_ctrl_10 == "democrats")TRUE -0.00941 0.00758
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:I(veto_1_ctrl_10 == "republicans")TRUE 0.00350 0.00570
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:court_review_10no 0.00034 0.00546
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:court_review_10yes 0.03650 0.01929
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:I(court_ctrl_10 == "democrats")TRUE 0.00127 0.00841
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:I(court_ctrl_10 == "republicans")TRUE -0.00117 0.00644
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:drawer_chg -0.00061 0.00843
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:drawer_ctrl_chg 0.00262 0.00603
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:veto_1_chg 0.00455 0.01008
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:veto_1_ctrl_chg -0.00175 0.00622
3c+ D.2 Abs. realized Nonpartisan Abs. efficiency gap, diff. from sims. dose_absparty:court_ctrl_chg -0.00526 0.00587
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst -0.50400 0.32490
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. nash_worst_10 0.05110 0.07834
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dvs_08 -0.05153 0.63250
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. is_south -0.25710 0.29000
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. log(n_distr_20) -0.27640 0.28390
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. distr_chg -0.02072 0.27080
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. log(1 + corrupt) 0.24090 0.20640
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. initiative 0.12130 0.23330
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. drawer_ctrl_chg -0.32770 0.27170
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. court_ctrl_chg -0.10420 0.22610
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:nash_worst_10 -0.00986 0.03814
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:dvs_08 -0.07472 0.26530
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:is_south 0.14360 0.19160
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:log(n_distr_20) 0.01850 0.05864
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:distr_chg 0.18210 0.16800
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:log(1 + corrupt) 0.00025 0.04151
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:initiative -0.13620 0.15710
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:drawer_ctrl_chg 0.13190 0.14440
4a Main Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:court_ctrl_chg 0.18680 0.17220
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst -0.18120 0.41060
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. nash_worst_10 -0.09741 0.15060
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dvs_08 0.53520 1.39800
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. is_south -0.37330 0.30410
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. log(n_distr_20) -0.30330 0.31690
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. distr_chg -0.12390 0.27530
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. log(1 + corrupt) 0.40550 0.21150
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. initiative -0.15480 0.25730
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. drawer_10commission -0.53550 0.72230
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. drawer_10legislature -0.01050 0.85270
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.55020 0.43560
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.05109 0.42290
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. I(veto_1_10 == "governor")TRUE -0.34140 0.57660
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.21420 0.30900
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE 0.57800 0.36340
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. court_review_10no -0.15960 0.30260
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. court_review_10yes 0.18410 0.39210
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. I(court_ctrl_10 == "democrats")TRUE -0.10190 0.32860
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.23580 0.30440
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. drawer_chg -0.28410 1.05400
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. drawer_ctrl_chg 0.01042 0.36960
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. veto_1_chg 0.70070 1.01800
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. veto_1_ctrl_chg -0.00776 0.32780
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. court_ctrl_chg -0.06676 0.27120
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:nash_worst_10 -0.00148 0.04059
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:dvs_08 -0.03020 0.26900
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:is_south 0.09836 0.21190
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:log(n_distr_20) 0.00439 0.06068
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:distr_chg 0.05349 0.17190
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:log(1 + corrupt) -0.00049 0.04250
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:initiative -0.12140 0.18440
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:drawer_10legislature -0.01497 0.14820
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:I(drawer_ctrl_10 == "democrats")TRUE -0.90500 0.43450
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:I(drawer_ctrl_10 == "republicans")TRUE 0.04001 0.15490
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:I(veto_1_10 == "governor")TRUE -0.02921 0.14830
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:I(veto_1_ctrl_10 == "democrats")TRUE 0.10410 0.27910
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:I(veto_1_ctrl_10 == "republicans")TRUE -0.05682 0.15820
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:court_review_10no 0.03578 0.15960
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:court_review_10yes 0.57100 0.44930
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:I(court_ctrl_10 == "democrats")TRUE -0.17900 0.32390
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:I(court_ctrl_10 == "republicans")TRUE -0.01615 0.17120
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:drawer_chg 0.07069 0.15410
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:drawer_ctrl_chg -0.00308 0.14690
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:veto_1_chg 0.07790 0.20590
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:veto_1_ctrl_chg -0.00425 0.19040
4a+ D.2 Maximum Nonpartisan Responsiveness, diff. from sims. dose_worst:court_ctrl_chg 0.07962 0.17550
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party -0.08573 0.13800
4b Main Realized Nonpartisan Responsiveness, diff. from sims. nash_party_10 -0.08401 0.06322
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dvs_08 0.09554 0.66180
4b Main Realized Nonpartisan Responsiveness, diff. from sims. is_south -0.04899 0.32920
4b Main Realized Nonpartisan Responsiveness, diff. from sims. log(n_distr_20) -0.35750 0.32270
4b Main Realized Nonpartisan Responsiveness, diff. from sims. distr_chg -0.12660 0.29450
4b Main Realized Nonpartisan Responsiveness, diff. from sims. log(1 + corrupt) 0.32660 0.22120
4b Main Realized Nonpartisan Responsiveness, diff. from sims. initiative 0.13080 0.24370
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party:nash_party_10 -0.00185 0.01646
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party:dvs_08 -0.00651 0.14760
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party:is_south 0.00086 0.07392
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party:log(n_distr_20) -0.00788 0.03970
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party:distr_chg 0.02144 0.07982
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party:log(1 + corrupt) -0.00627 0.02494
4b Main Realized Nonpartisan Responsiveness, diff. from sims. dose_party:initiative -0.04293 0.08384
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party -0.00599 0.17530
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. nash_party_10 -0.05105 0.12600
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dvs_08 -0.18630 1.56300
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. is_south -0.36350 0.38790
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. log(n_distr_20) -0.25150 0.34180
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. distr_chg -0.15370 0.34150
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. log(1 + corrupt) 0.37330 0.23640
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. initiative -0.17130 0.33910
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. drawer_10commission -0.14300 0.83420
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. drawer_10legislature -0.11400 0.86520
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.79430 0.62600
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.18270 0.55670
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. I(veto_1_10 == "governor")TRUE -0.41100 0.65980
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE 0.06975 0.43650
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE 0.76510 0.41020
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. court_review_10no -0.41080 0.34670
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. court_review_10yes -0.01165 0.43010
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.24440 0.40140
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.12310 0.38880
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. drawer_chg 0.27950 0.94210
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. drawer_ctrl_chg 0.28040 0.47880
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. veto_1_chg -0.00081 0.92100
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. veto_1_ctrl_chg 0.54730 0.39200
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. court_ctrl_chg -0.25840 0.29460
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:nash_party_10 0.00534 0.01986
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:dvs_08 0.01411 0.13920
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:is_south -0.00560 0.07863
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:log(n_distr_20) -0.00179 0.04120
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:distr_chg 0.00477 0.08758
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:log(1 + corrupt) -0.00097 0.02788
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:initiative -0.04540 0.09121
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:drawer_10legislature -0.00768 0.06933
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:I(drawer_ctrl_10 == "democrats")TRUE -0.02695 0.08248
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:I(drawer_ctrl_10 == "republicans")TRUE 0.03409 0.11240
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:I(veto_1_10 == "governor")TRUE -0.00871 0.06898
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:I(veto_1_ctrl_10 == "democrats")TRUE -0.06499 0.10610
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:I(veto_1_ctrl_10 == "republicans")TRUE 0.01809 0.10880
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:court_review_10no 0.04373 0.07101
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:court_review_10yes 0.03685 0.39190
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:I(court_ctrl_10 == "democrats")TRUE -0.02947 0.08249
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:I(court_ctrl_10 == "republicans")TRUE -0.00831 0.13180
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:drawer_chg 0.02206 0.17860
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:drawer_ctrl_chg -0.01627 0.06951
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:veto_1_chg 0.05711 0.21680
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:veto_1_ctrl_chg -0.00922 0.08647
4b+ D.2 Realized Nonpartisan Responsiveness, diff. from sims. dose_party:court_ctrl_chg 0.00783 0.07021
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty -0.00560 0.18900
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. nash_absparty_10 -0.00622 0.08726
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dvs_08 0.14080 0.70500
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. is_south -0.31190 0.33950
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. log(n_distr_20) -0.32170 0.30090
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. distr_chg -0.04463 0.28710
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. log(1 + corrupt) 0.28290 0.22800
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. initiative 0.11550 0.25450
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:nash_absparty_10 -0.00387 0.04031
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:dvs_08 0.02653 0.18060
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:is_south -0.03406 0.12830
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:log(n_distr_20) -0.00683 0.04464
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:distr_chg 0.00335 0.14030
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:log(1 + corrupt) -0.01062 0.03293
4c Main Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:initiative 0.03633 0.11470
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty -0.02750 0.25730
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. nash_absparty_10 -0.18790 0.13580
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dvs_08 0.39470 1.53400
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. is_south -0.46410 0.39350
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. log(n_distr_20) -0.31070 0.37590
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. distr_chg -0.10050 0.33720
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. log(1 + corrupt) 0.38090 0.26150
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. initiative -0.17360 0.34410
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. drawer_10commission -0.12580 0.81680
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. drawer_10legislature -0.04973 0.87700
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 1.23400 0.68870
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.37850 0.61120
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. I(veto_1_10 == "governor")TRUE -0.30140 0.73640
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE -0.12750 0.40940
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE 0.85990 0.41470
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. court_review_10no -0.38620 0.34690
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. court_review_10yes -0.13840 0.43670
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.12440 0.41160
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. I(court_ctrl_10 == "republicans")TRUE -0.17100 0.41870
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. drawer_chg 0.08526 0.89350
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. drawer_ctrl_chg 0.10690 0.48060
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. veto_1_chg 0.47600 0.96910
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. veto_1_ctrl_chg 0.22680 0.40200
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. court_ctrl_chg -0.26670 0.30540
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:nash_absparty_10 0.00089 0.04403
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:dvs_08 -0.00927 0.18100
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:is_south 0.03139 0.14320
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:log(n_distr_20) -0.00104 0.04650
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:distr_chg 0.01372 0.15880
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:log(1 + corrupt) -0.00044 0.03385
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:initiative -0.05271 0.12970
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:drawer_10legislature 0.00214 0.09770
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:I(drawer_ctrl_10 == "democrats")TRUE 0.01167 0.24690
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:I(drawer_ctrl_10 == "republicans")TRUE 0.00332 0.13090
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:I(veto_1_10 == "governor")TRUE -0.00298 0.09874
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:I(veto_1_ctrl_10 == "democrats")TRUE -0.00843 0.15740
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:I(veto_1_ctrl_10 == "republicans")TRUE -0.00023 0.12140
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:court_review_10no -0.02856 0.10910
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:court_review_10yes 0.21850 0.40620
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:I(court_ctrl_10 == "democrats")TRUE 0.08535 0.17580
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:I(court_ctrl_10 == "republicans")TRUE 0.01913 0.13920
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:drawer_chg 0.06450 0.16570
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:drawer_ctrl_chg -0.01714 0.12080
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:veto_1_chg 0.07506 0.21370
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:veto_1_ctrl_chg 0.03041 0.13440
4c+ D.2 Abs. realized Nonpartisan Responsiveness, diff. from sims. dose_absparty:court_ctrl_chg -0.01325 0.11660
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst -0.01655 0.05104
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. nash_worst_10 -0.00258 0.01322
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dvs_08 -0.06170 0.10770
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. is_south -0.05469 0.05171
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. log(n_distr_20) -0.01851 0.04371
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. distr_chg 0.00329 0.04656
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. log(1 + corrupt) 0.02261 0.03280
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. initiative -0.03212 0.03820
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. drawer_ctrl_chg 0.02625 0.04709
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. court_ctrl_chg 0.05630 0.03890
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:nash_worst_10 0.00083 0.00628
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:dvs_08 0.00381 0.03773
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:is_south -0.00756 0.03056
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:log(n_distr_20) 0.00028 0.00825
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:distr_chg -0.00109 0.02405
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:log(1 + corrupt) -0.00009 0.00652
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:initiative -0.00547 0.02586
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:drawer_ctrl_chg 0.00391 0.02188
5a Main Maximum Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_worst:court_ctrl_chg 0.01239 0.02484
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party 0.00630 0.01874
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. nash_party_10 -0.00272 0.00914
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dvs_08 -0.08287 0.09717
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. is_south -0.06912 0.04896
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. log(n_distr_20) 0.01363 0.04682
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. distr_chg -0.01491 0.04307
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. log(1 + corrupt) 0.01253 0.03231
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. initiative -0.02960 0.03439
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party:nash_party_10 -0.00315 0.00254
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party:dvs_08 0.00015 0.02003
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party:is_south 0.00365 0.01004
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party:log(n_distr_20) -0.00237 0.00564
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party:distr_chg 0.00227 0.01116
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party:log(1 + corrupt) -0.00032 0.00362
5b Main Realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_party:initiative -0.00019 0.01271
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty -0.00702 0.02875
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. nash_absparty_10 0.00090 0.01270
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dvs_08 -0.02874 0.10030
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. is_south -0.04410 0.04854
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. log(n_distr_20) -0.02017 0.04363
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. distr_chg 0.00346 0.04203
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. log(1 + corrupt) 0.01940 0.03195
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. initiative -0.02466 0.03779
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty:nash_absparty_10 -0.00130 0.00571
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty:dvs_08 -0.00227 0.02535
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty:is_south 0.00487 0.01834
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty:log(n_distr_20) -0.00142 0.00630
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty:distr_chg -0.00938 0.02096
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty:log(1 + corrupt) -0.00047 0.00447
5c Main Abs. realized Nonpartisan Abs. dilution asymmetry, diff. from sims. dose_absparty:initiative 0.00499 0.01662
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst 0.00602 0.01234
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. nash_worst_10 0.00402 0.00285
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dvs_08 0.01690 0.02290
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. is_south 0.00984 0.01086
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. log(n_distr_20) -0.00213 0.00934
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. distr_chg 0.03318 0.00950
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. log(1 + corrupt) -0.00940 0.00677
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. initiative 0.01186 0.00855
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. drawer_ctrl_chg -0.02433 0.01051
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. court_ctrl_chg 0.00093 0.00828
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:nash_worst_10 -0.00036 0.00154
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:dvs_08 -0.00173 0.01032
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:is_south 0.00573 0.00774
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:log(n_distr_20) 0.00024 0.00230
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:distr_chg 0.00054 0.00679
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:log(1 + corrupt) 0.00059 0.00164
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:initiative 0.00124 0.00610
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:drawer_ctrl_chg 0.00188 0.00548
6a Main Maximum Nonpartisan Abs. partisan bias, diff. from sims. dose_worst:court_ctrl_chg -0.00924 0.00629
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party -0.00203 0.00511
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. nash_party_10 -0.00297 0.00232
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dvs_08 0.02758 0.02401
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. is_south 0.00927 0.01205
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. log(n_distr_20) 0.00128 0.01184
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. distr_chg 0.02770 0.01077
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. log(1 + corrupt) -0.00801 0.00811
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. initiative -0.00106 0.00880
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party:nash_party_10 0.00100 0.00067
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party:dvs_08 -0.00056 0.00592
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party:is_south 0.00101 0.00271
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party:log(n_distr_20) 0.00096 0.00159
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party:distr_chg -0.00189 0.00309
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party:log(1 + corrupt) 0.00073 0.00104
6b Main Realized Nonpartisan Abs. partisan bias, diff. from sims. dose_party:initiative -0.00069 0.00326
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty 0.00429 0.00725
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. nash_absparty_10 0.00326 0.00326
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dvs_08 0.00928 0.02464
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. is_south -0.00256 0.01133
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. log(n_distr_20) 0.00405 0.01074
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. distr_chg 0.02870 0.00993
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. log(1 + corrupt) -0.00885 0.00792
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. initiative -0.00019 0.00886
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty:nash_absparty_10 0.00087 0.00153
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty:dvs_08 0.00034 0.00726
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty:is_south 0.00077 0.00479
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty:log(n_distr_20) 0.00052 0.00177
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty:distr_chg -0.00014 0.00510
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty:log(1 + corrupt) 0.00054 0.00126
6c Main Abs. realized Nonpartisan Abs. partisan bias, diff. from sims. dose_absparty:initiative 0.00164 0.00435
7a Main Maximum Nonpartisan Total partisan harm dose_worst -0.00846 0.01283
7a Main Maximum Nonpartisan Total partisan harm nash_worst_10 0.00078 0.00340
7a Main Maximum Nonpartisan Total partisan harm dvs_08 0.02159 0.02702
7a Main Maximum Nonpartisan Total partisan harm is_south 0.01918 0.01320
7a Main Maximum Nonpartisan Total partisan harm log(n_distr_20) -0.00395 0.01115
7a Main Maximum Nonpartisan Total partisan harm distr_chg -0.00567 0.01117
7a Main Maximum Nonpartisan Total partisan harm log(1 + corrupt) -0.00574 0.00842
7a Main Maximum Nonpartisan Total partisan harm initiative 0.01428 0.01012
7a Main Maximum Nonpartisan Total partisan harm drawer_ctrl_chg -0.01577 0.01222
7a Main Maximum Nonpartisan Total partisan harm court_ctrl_chg -0.01800 0.01012
7a Main Maximum Nonpartisan Total partisan harm dose_worst:nash_worst_10 -0.00026 0.00163
7a Main Maximum Nonpartisan Total partisan harm dose_worst:dvs_08 -0.00210 0.01078
7a Main Maximum Nonpartisan Total partisan harm dose_worst:is_south 0.00626 0.00784
7a Main Maximum Nonpartisan Total partisan harm dose_worst:log(n_distr_20) 0.00005 0.00236
7a Main Maximum Nonpartisan Total partisan harm dose_worst:distr_chg 0.00231 0.00661
7a Main Maximum Nonpartisan Total partisan harm dose_worst:log(1 + corrupt) 0.00009 0.00174
7a Main Maximum Nonpartisan Total partisan harm dose_worst:initiative 0.00042 0.00681
7a Main Maximum Nonpartisan Total partisan harm dose_worst:drawer_ctrl_chg 0.00261 0.00611
7a Main Maximum Nonpartisan Total partisan harm dose_worst:court_ctrl_chg 0.00054 0.00657
7b Main Realized Nonpartisan Total partisan harm dose_party -0.00043 0.00541
7b Main Realized Nonpartisan Total partisan harm nash_party_10 0.00172 0.00263
7b Main Realized Nonpartisan Total partisan harm dvs_08 0.02465 0.02414
7b Main Realized Nonpartisan Total partisan harm is_south 0.01746 0.01324
7b Main Realized Nonpartisan Total partisan harm log(n_distr_20) -0.01357 0.01212
7b Main Realized Nonpartisan Total partisan harm distr_chg 0.00178 0.01192
7b Main Realized Nonpartisan Total partisan harm log(1 + corrupt) -0.00224 0.00825
7b Main Realized Nonpartisan Total partisan harm initiative 0.01248 0.00925
7b Main Realized Nonpartisan Total partisan harm dose_party:nash_party_10 0.00088 0.00069
7b Main Realized Nonpartisan Total partisan harm dose_party:dvs_08 0.00053 0.00561
7b Main Realized Nonpartisan Total partisan harm dose_party:is_south -0.00142 0.00283
7b Main Realized Nonpartisan Total partisan harm dose_party:log(n_distr_20) 0.00077 0.00175
7b Main Realized Nonpartisan Total partisan harm dose_party:distr_chg -0.00080 0.00317
7b Main Realized Nonpartisan Total partisan harm dose_party:log(1 + corrupt) -0.00005 0.00104
7b Main Realized Nonpartisan Total partisan harm dose_party:initiative -0.00060 0.00332
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty 0.00256 0.00787
7c Main Abs. realized Nonpartisan Total partisan harm nash_absparty_10 -0.00094 0.00353
7c Main Abs. realized Nonpartisan Total partisan harm dvs_08 0.01376 0.02664
7c Main Abs. realized Nonpartisan Total partisan harm is_south 0.01270 0.01216
7c Main Abs. realized Nonpartisan Total partisan harm log(n_distr_20) -0.00639 0.01164
7c Main Abs. realized Nonpartisan Total partisan harm distr_chg -0.00479 0.01133
7c Main Abs. realized Nonpartisan Total partisan harm log(1 + corrupt) -0.00276 0.00902
7c Main Abs. realized Nonpartisan Total partisan harm initiative 0.00982 0.00980
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty:nash_absparty_10 0.00111 0.00163
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty:dvs_08 0.00052 0.00736
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty:is_south -0.00358 0.00514
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty:log(n_distr_20) 0.00025 0.00193
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty:distr_chg -0.00020 0.00559
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty:log(1 + corrupt) -0.00010 0.00136
7c Main Abs. realized Nonpartisan Total partisan harm dose_absparty:initiative -0.00160 0.00454
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst 0.19700 0.25760
8a Main Maximum Partisan Rep. seat diff. from sims. nash_worst_10 0.00178 0.06084
8a Main Maximum Partisan Rep. seat diff. from sims. dvs_08 -0.24920 0.47190
8a Main Maximum Partisan Rep. seat diff. from sims. is_south 0.51750 0.22440
8a Main Maximum Partisan Rep. seat diff. from sims. log(n_distr_20) 0.29080 0.20550
8a Main Maximum Partisan Rep. seat diff. from sims. distr_chg 0.10260 0.21440
8a Main Maximum Partisan Rep. seat diff. from sims. log(1 + corrupt) -0.21320 0.15390
8a Main Maximum Partisan Rep. seat diff. from sims. initiative 0.09136 0.17810
8a Main Maximum Partisan Rep. seat diff. from sims. drawer_ctrl_chg -0.19140 0.23010
8a Main Maximum Partisan Rep. seat diff. from sims. court_ctrl_chg -0.05284 0.17240
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:nash_worst_10 0.00198 0.03101
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:dvs_08 -0.00782 0.21250
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:is_south 0.05132 0.15350
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:log(n_distr_20) 0.00957 0.04585
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:distr_chg 0.06027 0.12600
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:log(1 + corrupt) 0.00891 0.03472
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:initiative -0.01056 0.13280
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:drawer_ctrl_chg -0.06641 0.11400
8a Main Maximum Partisan Rep. seat diff. from sims. dose_worst:court_ctrl_chg 0.06847 0.12940
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst 0.20810 0.35170
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. nash_worst_10 0.02348 0.14290
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dvs_08 -0.76860 1.22600
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. is_south 0.56440 0.30830
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. log(n_distr_20) 0.30180 0.27890
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. distr_chg 0.14830 0.26970
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. log(1 + corrupt) -0.24640 0.19720
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. initiative 0.09653 0.25030
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. drawer_10commission 0.26620 0.65970
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. drawer_10legislature 0.01679 0.76260
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE -0.06916 0.42040
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE -0.01912 0.37730
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. I(veto_1_10 == "governor")TRUE 0.24340 0.51330
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE -0.26930 0.32690
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.26530 0.35180
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. court_review_10no 0.21690 0.31100
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. court_review_10yes 0.11130 0.39120
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.04582 0.32590
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. I(court_ctrl_10 == "republicans")TRUE 0.19380 0.33850
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. drawer_chg 0.27910 0.92880
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. drawer_ctrl_chg -0.30780 0.35500
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. veto_1_chg 0.08962 0.85470
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. veto_1_ctrl_chg -0.04079 0.31050
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. court_ctrl_chg -0.04083 0.25880
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:nash_worst_10 0.00188 0.03223
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:dvs_08 0.02282 0.21050
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:is_south 0.03612 0.17970
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:log(n_distr_20) 0.00886 0.04770
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:distr_chg 0.04877 0.13570
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:log(1 + corrupt) 0.00737 0.03647
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:initiative -0.02515 0.15770
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:drawer_10legislature 0.00944 0.13000
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:I(drawer_ctrl_10 == "democrats")TRUE -0.17910 0.37340
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:I(drawer_ctrl_10 == "republicans")TRUE 0.04617 0.11580
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:I(veto_1_10 == "governor")TRUE -0.00715 0.13200
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:I(veto_1_ctrl_10 == "democrats")TRUE -0.05867 0.24030
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:I(veto_1_ctrl_10 == "republicans")TRUE 0.00863 0.13090
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:court_review_10no -0.01471 0.12560
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:court_review_10yes -0.01978 0.38170
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:I(court_ctrl_10 == "democrats")TRUE -0.11270 0.29250
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:I(court_ctrl_10 == "republicans")TRUE 0.06086 0.14210
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:drawer_chg -0.02472 0.12740
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:drawer_ctrl_chg -0.02135 0.12300
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:veto_1_chg -0.03994 0.17040
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:veto_1_ctrl_chg 0.00901 0.16070
8a+ D.2 Maximum Partisan Rep. seat diff. from sims. dose_worst:court_ctrl_chg 0.05177 0.14100
8b Main Realized Partisan Rep. seat diff. from sims. dose_party 0.02553 0.10550
8b Main Realized Partisan Rep. seat diff. from sims. nash_party_10 0.04329 0.04348
8b Main Realized Partisan Rep. seat diff. from sims. dvs_08 -0.36990 0.42340
8b Main Realized Partisan Rep. seat diff. from sims. is_south 0.32160 0.21790
8b Main Realized Partisan Rep. seat diff. from sims. log(n_distr_20) 0.39780 0.21440
8b Main Realized Partisan Rep. seat diff. from sims. distr_chg 0.00060 0.19790
8b Main Realized Partisan Rep. seat diff. from sims. log(1 + corrupt) -0.26750 0.14720
8b Main Realized Partisan Rep. seat diff. from sims. initiative 0.06358 0.16670
8b Main Realized Partisan Rep. seat diff. from sims. dose_party:nash_party_10 0.02353 0.01387
8b Main Realized Partisan Rep. seat diff. from sims. dose_party:dvs_08 0.03293 0.11680
8b Main Realized Partisan Rep. seat diff. from sims. dose_party:is_south -0.02686 0.06378
8b Main Realized Partisan Rep. seat diff. from sims. dose_party:log(n_distr_20) 0.03512 0.03376
8b Main Realized Partisan Rep. seat diff. from sims. dose_party:distr_chg 0.08301 0.06018
8b Main Realized Partisan Rep. seat diff. from sims. dose_party:log(1 + corrupt) 0.01168 0.01962
8b Main Realized Partisan Rep. seat diff. from sims. dose_party:initiative -0.00112 0.06158
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party 0.01594 0.14230
8b+ D.2 Realized Partisan Rep. seat diff. from sims. nash_party_10 0.07948 0.11180
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dvs_08 -1.10000 1.33900
8b+ D.2 Realized Partisan Rep. seat diff. from sims. is_south 0.44200 0.32090
8b+ D.2 Realized Partisan Rep. seat diff. from sims. log(n_distr_20) 0.39540 0.29060
8b+ D.2 Realized Partisan Rep. seat diff. from sims. distr_chg 0.17260 0.28760
8b+ D.2 Realized Partisan Rep. seat diff. from sims. log(1 + corrupt) -0.30080 0.20260
8b+ D.2 Realized Partisan Rep. seat diff. from sims. initiative 0.00308 0.25690
8b+ D.2 Realized Partisan Rep. seat diff. from sims. drawer_10commission 0.47490 0.68780
8b+ D.2 Realized Partisan Rep. seat diff. from sims. drawer_10legislature 0.11430 0.69290
8b+ D.2 Realized Partisan Rep. seat diff. from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.15080 0.52610
8b+ D.2 Realized Partisan Rep. seat diff. from sims. I(drawer_ctrl_10 == "republicans")TRUE -0.22670 0.44990
8b+ D.2 Realized Partisan Rep. seat diff. from sims. I(veto_1_10 == "governor")TRUE 0.46980 0.49860
8b+ D.2 Realized Partisan Rep. seat diff. from sims. I(veto_1_ctrl_10 == "democrats")TRUE -0.29870 0.35920
8b+ D.2 Realized Partisan Rep. seat diff. from sims. I(veto_1_ctrl_10 == "republicans")TRUE -0.42880 0.34130
8b+ D.2 Realized Partisan Rep. seat diff. from sims. court_review_10no 0.30910 0.27590
8b+ D.2 Realized Partisan Rep. seat diff. from sims. court_review_10yes 0.19310 0.35550
8b+ D.2 Realized Partisan Rep. seat diff. from sims. I(court_ctrl_10 == "democrats")TRUE 0.04485 0.31810
8b+ D.2 Realized Partisan Rep. seat diff. from sims. I(court_ctrl_10 == "republicans")TRUE 0.07080 0.30440
8b+ D.2 Realized Partisan Rep. seat diff. from sims. drawer_chg -0.47210 0.77770
8b+ D.2 Realized Partisan Rep. seat diff. from sims. drawer_ctrl_chg -0.15930 0.39350
8b+ D.2 Realized Partisan Rep. seat diff. from sims. veto_1_chg 0.30750 0.77180
8b+ D.2 Realized Partisan Rep. seat diff. from sims. veto_1_ctrl_chg -0.00968 0.33480
8b+ D.2 Realized Partisan Rep. seat diff. from sims. court_ctrl_chg -0.09399 0.25160
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:nash_party_10 0.00834 0.01666
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:dvs_08 0.01338 0.11850
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:is_south -0.00680 0.06296
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:log(n_distr_20) 0.01532 0.03347
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:distr_chg 0.02749 0.07040
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:log(1 + corrupt) 0.00462 0.02128
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:initiative 0.00800 0.07413
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:drawer_10legislature 0.00351 0.05075
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:I(drawer_ctrl_10 == "democrats")TRUE -0.00862 0.06809
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:I(drawer_ctrl_10 == "republicans")TRUE 0.04945 0.09637
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:I(veto_1_10 == "governor")TRUE -0.01114 0.05533
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:I(veto_1_ctrl_10 == "democrats")TRUE 0.00202 0.08189
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:I(veto_1_ctrl_10 == "republicans")TRUE 0.00427 0.09334
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:court_review_10no -0.02857 0.06083
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:court_review_10yes 0.24700 0.33160
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:I(court_ctrl_10 == "democrats")TRUE -0.00641 0.06353
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:I(court_ctrl_10 == "republicans")TRUE 0.04808 0.11920
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:drawer_chg 0.03371 0.13870
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:drawer_ctrl_chg -0.01517 0.06202
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:veto_1_chg 0.01649 0.18190
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:veto_1_ctrl_chg 0.01954 0.07414
8b+ D.2 Realized Partisan Rep. seat diff. from sims. dose_party:court_ctrl_chg 0.00556 0.05652
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst 1.73700 1.42200
9a Main Maximum Partisan Rep. seats z-score from sims. nash_worst_10 -0.15800 0.35280
9a Main Maximum Partisan Rep. seats z-score from sims. dvs_08 0.09149 2.75900
9a Main Maximum Partisan Rep. seats z-score from sims. is_south 2.42000 1.33200
9a Main Maximum Partisan Rep. seats z-score from sims. log(n_distr_20) 1.17400 1.18400
9a Main Maximum Partisan Rep. seats z-score from sims. distr_chg -0.92050 1.15800
9a Main Maximum Partisan Rep. seats z-score from sims. log(1 + corrupt) -0.86700 0.88810
9a Main Maximum Partisan Rep. seats z-score from sims. initiative 0.14550 1.08000
9a Main Maximum Partisan Rep. seats z-score from sims. drawer_ctrl_chg 0.02843 1.33500
9a Main Maximum Partisan Rep. seats z-score from sims. court_ctrl_chg -0.07205 1.01300
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:nash_worst_10 0.05404 0.17690
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:dvs_08 0.08201 1.20400
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:is_south 0.14050 0.87270
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:log(n_distr_20) 0.01541 0.26660
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:distr_chg -0.20640 0.76030
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:log(1 + corrupt) 0.05175 0.19320
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:initiative 0.03157 0.73720
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:drawer_ctrl_chg -0.44930 0.63990
9a Main Maximum Partisan Rep. seats z-score from sims. dose_worst:court_ctrl_chg 0.17870 0.75770
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst 0.80630 1.92800
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. nash_worst_10 0.16740 0.83760
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dvs_08 -2.25000 7.65000
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. is_south 2.49700 1.65400
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. log(n_distr_20) 0.97050 1.68700
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. distr_chg -0.99860 1.62200
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. log(1 + corrupt) -0.87270 1.13700
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. initiative 0.44200 1.45800
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. drawer_10commission 1.15400 3.88400
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. drawer_10legislature 0.45130 4.18300
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. I(drawer_ctrl_10 == "democrats")TRUE -0.11380 2.51500
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. I(drawer_ctrl_10 == "republicans")TRUE 0.43320 2.17300
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. I(veto_1_10 == "governor")TRUE 0.38540 3.15800
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. I(veto_1_ctrl_10 == "democrats")TRUE -1.02600 1.96000
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. I(veto_1_ctrl_10 == "republicans")TRUE -1.45400 2.05800
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. court_review_10no 0.93910 1.86000
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. court_review_10yes 1.11300 2.18200
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. I(court_ctrl_10 == "democrats")TRUE 0.05338 1.80400
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. I(court_ctrl_10 == "republicans")TRUE 0.11520 1.82900
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. drawer_chg -0.75450 5.49100
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. drawer_ctrl_chg -0.75670 2.10600
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. veto_1_chg 1.49200 5.05100
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. veto_1_ctrl_chg -1.58700 1.90900
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. court_ctrl_chg 0.16180 1.54700
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:nash_worst_10 0.02366 0.18430
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:dvs_08 0.05374 1.25600
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:is_south 0.10750 1.01100
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:log(n_distr_20) 0.00620 0.25900
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:distr_chg 0.06184 0.74510
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:log(1 + corrupt) 0.01502 0.18940
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:initiative -0.09447 0.86890
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:drawer_10legislature 0.05209 0.67860
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:I(drawer_ctrl_10 == "democrats")TRUE -0.01665 2.06000
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:I(drawer_ctrl_10 == "republicans")TRUE 0.23780 0.72980
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:I(veto_1_10 == "governor")TRUE 0.02656 0.65680
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:I(veto_1_ctrl_10 == "democrats")TRUE -0.57960 1.36100
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:I(veto_1_ctrl_10 == "republicans")TRUE 0.20810 0.74810
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:court_review_10no -0.05030 0.74550
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:court_review_10yes -0.30720 2.10500
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:I(court_ctrl_10 == "democrats")TRUE -0.30670 1.63100
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:I(court_ctrl_10 == "republicans")TRUE 0.09971 0.83590
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:drawer_chg -0.05107 0.69360
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:drawer_ctrl_chg -0.12310 0.71940
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:veto_1_chg -0.26720 0.95170
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:veto_1_ctrl_chg 0.48540 0.97680
9a+ D.2 Maximum Partisan Rep. seats z-score from sims. dose_worst:court_ctrl_chg 0.26320 0.83240
9b Main Realized Partisan Rep. seats z-score from sims. dose_party 0.34900 0.57950
9b Main Realized Partisan Rep. seats z-score from sims. nash_party_10 0.28230 0.25220
9b Main Realized Partisan Rep. seats z-score from sims. dvs_08 -0.89330 2.52500
9b Main Realized Partisan Rep. seats z-score from sims. is_south 1.25300 1.24900
9b Main Realized Partisan Rep. seats z-score from sims. log(n_distr_20) 1.60500 1.27200
9b Main Realized Partisan Rep. seats z-score from sims. distr_chg -1.02600 1.19500
9b Main Realized Partisan Rep. seats z-score from sims. log(1 + corrupt) -1.28400 0.86210
9b Main Realized Partisan Rep. seats z-score from sims. initiative 0.19900 0.94050
9b Main Realized Partisan Rep. seats z-score from sims. dose_party:nash_party_10 0.12080 0.07735
9b Main Realized Partisan Rep. seats z-score from sims. dose_party:dvs_08 0.16350 0.61310
9b Main Realized Partisan Rep. seats z-score from sims. dose_party:is_south -0.14580 0.33140
9b Main Realized Partisan Rep. seats z-score from sims. dose_party:log(n_distr_20) 0.15040 0.18520
9b Main Realized Partisan Rep. seats z-score from sims. dose_party:distr_chg 0.25030 0.32810
9b Main Realized Partisan Rep. seats z-score from sims. dose_party:log(1 + corrupt) 0.05876 0.11610
9b Main Realized Partisan Rep. seats z-score from sims. dose_party:initiative -0.05912 0.37160
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party 0.08853 0.82570
9b+ D.2 Realized Partisan Rep. seats z-score from sims. nash_party_10 0.71830 0.62430
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dvs_08 -2.04100 7.07200
9b+ D.2 Realized Partisan Rep. seats z-score from sims. is_south 2.15400 1.76500
9b+ D.2 Realized Partisan Rep. seats z-score from sims. log(n_distr_20) 1.11800 1.69200
9b+ D.2 Realized Partisan Rep. seats z-score from sims. distr_chg -0.69270 1.58100
9b+ D.2 Realized Partisan Rep. seats z-score from sims. log(1 + corrupt) -0.98040 1.11500
9b+ D.2 Realized Partisan Rep. seats z-score from sims. initiative 0.08137 1.58900
9b+ D.2 Realized Partisan Rep. seats z-score from sims. drawer_10commission 1.13200 3.78300
9b+ D.2 Realized Partisan Rep. seats z-score from sims. drawer_10legislature 0.70870 3.87500
9b+ D.2 Realized Partisan Rep. seats z-score from sims. I(drawer_ctrl_10 == "democrats")TRUE 0.82820 3.04500
9b+ D.2 Realized Partisan Rep. seats z-score from sims. I(drawer_ctrl_10 == "republicans")TRUE -1.72200 2.71900
9b+ D.2 Realized Partisan Rep. seats z-score from sims. I(veto_1_10 == "governor")TRUE 1.44800 2.97200
9b+ D.2 Realized Partisan Rep. seats z-score from sims. I(veto_1_ctrl_10 == "democrats")TRUE -0.80700 2.16100
9b+ D.2 Realized Partisan Rep. seats z-score from sims. I(veto_1_ctrl_10 == "republicans")TRUE -2.31200 1.89100
9b+ D.2 Realized Partisan Rep. seats z-score from sims. court_review_10no 1.32000 1.56000
9b+ D.2 Realized Partisan Rep. seats z-score from sims. court_review_10yes 1.55200 2.02700
9b+ D.2 Realized Partisan Rep. seats z-score from sims. I(court_ctrl_10 == "democrats")TRUE -0.71870 1.92900
9b+ D.2 Realized Partisan Rep. seats z-score from sims. I(court_ctrl_10 == "republicans")TRUE -0.55910 1.86500
9b+ D.2 Realized Partisan Rep. seats z-score from sims. drawer_chg -3.72200 4.07000
9b+ D.2 Realized Partisan Rep. seats z-score from sims. drawer_ctrl_chg -0.13950 2.36800
9b+ D.2 Realized Partisan Rep. seats z-score from sims. veto_1_chg 2.51600 4.20700
9b+ D.2 Realized Partisan Rep. seats z-score from sims. veto_1_ctrl_chg -2.19100 1.86100
9b+ D.2 Realized Partisan Rep. seats z-score from sims. court_ctrl_chg 0.33620 1.38200
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:nash_party_10 0.02655 0.10370
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:dvs_08 0.00483 0.62540
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:is_south -0.04476 0.33730
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:log(n_distr_20) 0.04511 0.19360
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:distr_chg 0.06974 0.42770
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:log(1 + corrupt) 0.01071 0.11620
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:initiative -0.05276 0.41630
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:drawer_10legislature 0.02427 0.31970
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:I(drawer_ctrl_10 == "democrats")TRUE -0.03177 0.40130
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:I(drawer_ctrl_10 == "republicans")TRUE 0.28500 0.53530
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:I(veto_1_10 == "governor")TRUE -0.02319 0.30400
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:I(veto_1_ctrl_10 == "democrats")TRUE 0.06641 0.47490
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:I(veto_1_ctrl_10 == "republicans")TRUE -0.00365 0.55150
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:court_review_10no -0.15950 0.34320
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:court_review_10yes 1.08200 1.79200
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:I(court_ctrl_10 == "democrats")TRUE 0.02149 0.37710
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:I(court_ctrl_10 == "republicans")TRUE 0.19370 0.62550
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:drawer_chg 0.22380 0.81840
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:drawer_ctrl_chg -0.06979 0.33330
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:veto_1_chg 0.14020 1.02000
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:veto_1_ctrl_chg 0.21440 0.41410
9b+ D.2 Realized Partisan Rep. seats z-score from sims. dose_party:court_ctrl_chg 0.07013 0.30920
10a Main Maximum Partisan Efficiency gap dose_worst 0.01197 0.02949
10a Main Maximum Partisan Efficiency gap nash_worst_10 0.00648 0.00778
10a Main Maximum Partisan Efficiency gap dvs_08 -0.01810 0.06550
10a Main Maximum Partisan Efficiency gap is_south -0.02360 0.02893
10a Main Maximum Partisan Efficiency gap log(n_distr_20) 0.02856 0.02624
10a Main Maximum Partisan Efficiency gap distr_chg -0.02886 0.02671
10a Main Maximum Partisan Efficiency gap log(1 + corrupt) -0.01623 0.01974
10a Main Maximum Partisan Efficiency gap initiative -0.00091 0.02398
10a Main Maximum Partisan Efficiency gap drawer_ctrl_chg -0.01770 0.02892
10a Main Maximum Partisan Efficiency gap court_ctrl_chg -0.00948 0.02454
10a Main Maximum Partisan Efficiency gap dose_worst:nash_worst_10 0.00064 0.00364
10a Main Maximum Partisan Efficiency gap dose_worst:dvs_08 0.00053 0.02227
10a Main Maximum Partisan Efficiency gap dose_worst:is_south 0.00553 0.01810
10a Main Maximum Partisan Efficiency gap dose_worst:log(n_distr_20) -0.00020 0.00530
10a Main Maximum Partisan Efficiency gap dose_worst:distr_chg -0.00541 0.01536
10a Main Maximum Partisan Efficiency gap dose_worst:log(1 + corrupt) 0.00018 0.00384
10a Main Maximum Partisan Efficiency gap dose_worst:initiative -0.00338 0.01460
10a Main Maximum Partisan Efficiency gap dose_worst:drawer_ctrl_chg -0.00540 0.01300
10a Main Maximum Partisan Efficiency gap dose_worst:court_ctrl_chg -0.00076 0.01453
10a+ D.2 Maximum Partisan Efficiency gap dose_worst 0.01498 0.03745
10a+ D.2 Maximum Partisan Efficiency gap nash_worst_10 0.01136 0.01734
10a+ D.2 Maximum Partisan Efficiency gap dvs_08 0.03839 0.15410
10a+ D.2 Maximum Partisan Efficiency gap is_south -0.00930 0.03735
10a+ D.2 Maximum Partisan Efficiency gap log(n_distr_20) 0.01005 0.03332
10a+ D.2 Maximum Partisan Efficiency gap distr_chg -0.02365 0.03164
10a+ D.2 Maximum Partisan Efficiency gap log(1 + corrupt) -0.00185 0.02359
10a+ D.2 Maximum Partisan Efficiency gap initiative 0.02023 0.03262
10a+ D.2 Maximum Partisan Efficiency gap drawer_10commission -0.03986 0.07601
10a+ D.2 Maximum Partisan Efficiency gap drawer_10legislature -0.04663 0.08777
10a+ D.2 Maximum Partisan Efficiency gap I(drawer_ctrl_10 == "democrats")TRUE 0.00298 0.05292
10a+ D.2 Maximum Partisan Efficiency gap I(drawer_ctrl_10 == "republicans")TRUE 0.01184 0.04627
10a+ D.2 Maximum Partisan Efficiency gap I(veto_1_10 == "governor")TRUE -0.04806 0.06492
10a+ D.2 Maximum Partisan Efficiency gap I(veto_1_ctrl_10 == "democrats")TRUE 0.01952 0.03798
10a+ D.2 Maximum Partisan Efficiency gap I(veto_1_ctrl_10 == "republicans")TRUE 0.02862 0.04525
10a+ D.2 Maximum Partisan Efficiency gap court_review_10no 0.01190 0.03807
10a+ D.2 Maximum Partisan Efficiency gap court_review_10yes 0.01399 0.05064
10a+ D.2 Maximum Partisan Efficiency gap I(court_ctrl_10 == "democrats")TRUE -0.01871 0.04009
10a+ D.2 Maximum Partisan Efficiency gap I(court_ctrl_10 == "republicans")TRUE -0.00957 0.04003
10a+ D.2 Maximum Partisan Efficiency gap drawer_chg 0.06741 0.10920
10a+ D.2 Maximum Partisan Efficiency gap drawer_ctrl_chg -0.02960 0.04368
10a+ D.2 Maximum Partisan Efficiency gap veto_1_chg 0.04436 0.10240
10a+ D.2 Maximum Partisan Efficiency gap veto_1_ctrl_chg 0.00455 0.04001
10a+ D.2 Maximum Partisan Efficiency gap court_ctrl_chg 0.00182 0.03145
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:nash_worst_10 0.00078 0.00394
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:dvs_08 0.00188 0.02419
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:is_south 0.00688 0.01959
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:log(n_distr_20) 0.00080 0.00513
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:distr_chg -0.00246 0.01596
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:log(1 + corrupt) 0.00027 0.00408
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:initiative -0.00479 0.01704
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:drawer_10legislature 0.00068 0.01253
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:I(drawer_ctrl_10 == "democrats")TRUE 0.00509 0.04035
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:I(drawer_ctrl_10 == "republicans")TRUE 0.00086 0.01407
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:I(veto_1_10 == "governor")TRUE 0.00116 0.01411
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:I(veto_1_ctrl_10 == "democrats")TRUE -0.00337 0.02816
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:I(veto_1_ctrl_10 == "republicans")TRUE 0.00154 0.01479
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:court_review_10no 0.00178 0.01435
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:court_review_10yes -0.00774 0.04338
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:I(court_ctrl_10 == "democrats")TRUE -0.00238 0.03144
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:I(court_ctrl_10 == "republicans")TRUE -0.00067 0.01611
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:drawer_chg 0.00221 0.01422
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:drawer_ctrl_chg 0.00033 0.01400
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:veto_1_chg 0.00412 0.01880
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:veto_1_ctrl_chg 0.00869 0.01927
10a+ D.2 Maximum Partisan Efficiency gap dose_worst:court_ctrl_chg 0.00024 0.01612
10b Main Realized Partisan Efficiency gap dose_party -0.00743 0.01247
10b Main Realized Partisan Efficiency gap nash_party_10 0.00543 0.00563
10b Main Realized Partisan Efficiency gap dvs_08 0.02552 0.05499
10b Main Realized Partisan Efficiency gap is_south -0.00962 0.02864
10b Main Realized Partisan Efficiency gap log(n_distr_20) 0.00654 0.02732
10b Main Realized Partisan Efficiency gap distr_chg -0.01610 0.02644
10b Main Realized Partisan Efficiency gap log(1 + corrupt) -0.01025 0.01833
10b Main Realized Partisan Efficiency gap initiative 0.00352 0.02253
10b Main Realized Partisan Efficiency gap dose_party:nash_party_10 0.00101 0.00158
10b Main Realized Partisan Efficiency gap dose_party:dvs_08 0.00011 0.01222
10b Main Realized Partisan Efficiency gap dose_party:is_south -0.00039 0.00688
10b Main Realized Partisan Efficiency gap dose_party:log(n_distr_20) 0.00159 0.00361
10b Main Realized Partisan Efficiency gap dose_party:distr_chg -0.00111 0.00709
10b Main Realized Partisan Efficiency gap dose_party:log(1 + corrupt) 0.00053 0.00220
10b Main Realized Partisan Efficiency gap dose_party:initiative 0.00214 0.00754
10b+ D.2 Realized Partisan Efficiency gap dose_party -0.00775 0.01615
10b+ D.2 Realized Partisan Efficiency gap nash_party_10 0.01562 0.01267
10b+ D.2 Realized Partisan Efficiency gap dvs_08 0.00046 0.15650
10b+ D.2 Realized Partisan Efficiency gap is_south 0.00778 0.03951
10b+ D.2 Realized Partisan Efficiency gap log(n_distr_20) 0.01150 0.03579
10b+ D.2 Realized Partisan Efficiency gap distr_chg -0.01931 0.03319
10b+ D.2 Realized Partisan Efficiency gap log(1 + corrupt) -0.00591 0.02551
10b+ D.2 Realized Partisan Efficiency gap initiative 0.00916 0.03327
10b+ D.2 Realized Partisan Efficiency gap drawer_10commission -0.01549 0.08180
10b+ D.2 Realized Partisan Efficiency gap drawer_10legislature -0.02480 0.08173
10b+ D.2 Realized Partisan Efficiency gap I(drawer_ctrl_10 == "democrats")TRUE 0.04688 0.06259
10b+ D.2 Realized Partisan Efficiency gap I(drawer_ctrl_10 == "republicans")TRUE -0.02916 0.05312
10b+ D.2 Realized Partisan Efficiency gap I(veto_1_10 == "governor")TRUE -0.00717 0.06221
10b+ D.2 Realized Partisan Efficiency gap I(veto_1_ctrl_10 == "democrats")TRUE 0.02335 0.04452
10b+ D.2 Realized Partisan Efficiency gap I(veto_1_ctrl_10 == "republicans")TRUE -0.00783 0.04256
10b+ D.2 Realized Partisan Efficiency gap court_review_10no 0.03913 0.03538
10b+ D.2 Realized Partisan Efficiency gap court_review_10yes 0.01345 0.04314
10b+ D.2 Realized Partisan Efficiency gap I(court_ctrl_10 == "democrats")TRUE -0.01611 0.03973
10b+ D.2 Realized Partisan Efficiency gap I(court_ctrl_10 == "republicans")TRUE -0.01268 0.04126
10b+ D.2 Realized Partisan Efficiency gap drawer_chg 0.00320 0.08687
10b+ D.2 Realized Partisan Efficiency gap drawer_ctrl_chg -0.00341 0.05033
10b+ D.2 Realized Partisan Efficiency gap veto_1_chg 0.02487 0.09116
10b+ D.2 Realized Partisan Efficiency gap veto_1_ctrl_chg -0.00374 0.04044
10b+ D.2 Realized Partisan Efficiency gap court_ctrl_chg 0.00655 0.02976
10b+ D.2 Realized Partisan Efficiency gap dose_party:nash_party_10 0.00027 0.00188
10b+ D.2 Realized Partisan Efficiency gap dose_party:dvs_08 -0.00067 0.01289
10b+ D.2 Realized Partisan Efficiency gap dose_party:is_south -0.00037 0.00702
10b+ D.2 Realized Partisan Efficiency gap dose_party:log(n_distr_20) 0.00014 0.00386
10b+ D.2 Realized Partisan Efficiency gap dose_party:distr_chg -0.00141 0.00784
10b+ D.2 Realized Partisan Efficiency gap dose_party:log(1 + corrupt) -0.00034 0.00225
10b+ D.2 Realized Partisan Efficiency gap dose_party:initiative 0.00058 0.00882
10b+ D.2 Realized Partisan Efficiency gap dose_party:drawer_10legislature -0.00011 0.00612
10b+ D.2 Realized Partisan Efficiency gap dose_party:I(drawer_ctrl_10 == "democrats")TRUE 0.00033 0.00794
10b+ D.2 Realized Partisan Efficiency gap dose_party:I(drawer_ctrl_10 == "republicans")TRUE 0.00241 0.01110
10b+ D.2 Realized Partisan Efficiency gap dose_party:I(veto_1_10 == "governor")TRUE -0.00049 0.00582
10b+ D.2 Realized Partisan Efficiency gap dose_party:I(veto_1_ctrl_10 == "democrats")TRUE 0.00077 0.00994
10b+ D.2 Realized Partisan Efficiency gap dose_party:I(veto_1_ctrl_10 == "republicans")TRUE 0.00268 0.01104
10b+ D.2 Realized Partisan Efficiency gap dose_party:court_review_10no -0.00153 0.00658
10b+ D.2 Realized Partisan Efficiency gap dose_party:court_review_10yes 0.02675 0.03923
10b+ D.2 Realized Partisan Efficiency gap dose_party:I(court_ctrl_10 == "democrats")TRUE -0.00212 0.00775
10b+ D.2 Realized Partisan Efficiency gap dose_party:I(court_ctrl_10 == "republicans")TRUE 0.00294 0.01269
10b+ D.2 Realized Partisan Efficiency gap dose_party:drawer_chg 0.00908 0.01552
10b+ D.2 Realized Partisan Efficiency gap dose_party:drawer_ctrl_chg -0.00172 0.00663
10b+ D.2 Realized Partisan Efficiency gap dose_party:veto_1_chg 0.01886 0.02114
10b+ D.2 Realized Partisan Efficiency gap dose_party:veto_1_ctrl_chg 0.00521 0.00885
10b+ D.2 Realized Partisan Efficiency gap dose_party:court_ctrl_chg 0.00030 0.00639
11a Main Maximum Partisan Dilution asymmetry dose_worst 0.04456 0.05713
11a Main Maximum Partisan Dilution asymmetry nash_worst_10 0.02352 0.01541
11a Main Maximum Partisan Dilution asymmetry dvs_08 0.17540 0.12330
11a Main Maximum Partisan Dilution asymmetry is_south -0.04267 0.05469
11a Main Maximum Partisan Dilution asymmetry log(n_distr_20) -0.03926 0.05085
11a Main Maximum Partisan Dilution asymmetry distr_chg -0.01374 0.05356
11a Main Maximum Partisan Dilution asymmetry log(1 + corrupt) -0.00361 0.03812
11a Main Maximum Partisan Dilution asymmetry initiative -0.04141 0.04793
11a Main Maximum Partisan Dilution asymmetry drawer_ctrl_chg -0.02521 0.05771
11a Main Maximum Partisan Dilution asymmetry court_ctrl_chg -0.01514 0.04575
11a Main Maximum Partisan Dilution asymmetry dose_worst:nash_worst_10 0.00088 0.00723
11a Main Maximum Partisan Dilution asymmetry dose_worst:dvs_08 0.00068 0.04689
11a Main Maximum Partisan Dilution asymmetry dose_worst:is_south 0.02063 0.03611
11a Main Maximum Partisan Dilution asymmetry dose_worst:log(n_distr_20) -0.00164 0.01035
11a Main Maximum Partisan Dilution asymmetry dose_worst:distr_chg -0.01983 0.03073
11a Main Maximum Partisan Dilution asymmetry dose_worst:log(1 + corrupt) -0.00004 0.00795
11a Main Maximum Partisan Dilution asymmetry dose_worst:initiative -0.01498 0.03242
11a Main Maximum Partisan Dilution asymmetry dose_worst:drawer_ctrl_chg -0.01468 0.02716
11a Main Maximum Partisan Dilution asymmetry dose_worst:court_ctrl_chg -0.00772 0.02914
11b Main Realized Partisan Dilution asymmetry dose_party -0.01136 0.02396
11b Main Realized Partisan Dilution asymmetry nash_party_10 0.01918 0.01088
11b Main Realized Partisan Dilution asymmetry dvs_08 0.25380 0.11200
11b Main Realized Partisan Dilution asymmetry is_south -0.04820 0.05763
11b Main Realized Partisan Dilution asymmetry log(n_distr_20) -0.06831 0.05336
11b Main Realized Partisan Dilution asymmetry distr_chg 0.00387 0.05346
11b Main Realized Partisan Dilution asymmetry log(1 + corrupt) 0.00498 0.03690
11b Main Realized Partisan Dilution asymmetry initiative -0.02508 0.04495
11b Main Realized Partisan Dilution asymmetry dose_party:nash_party_10 -0.00084 0.00319
11b Main Realized Partisan Dilution asymmetry dose_party:dvs_08 -0.00502 0.02511
11b Main Realized Partisan Dilution asymmetry dose_party:is_south 0.01132 0.01358
11b Main Realized Partisan Dilution asymmetry dose_party:log(n_distr_20) 0.00247 0.00729
11b Main Realized Partisan Dilution asymmetry dose_party:distr_chg 0.00022 0.01351
11b Main Realized Partisan Dilution asymmetry dose_party:log(1 + corrupt) 0.00133 0.00463
11b Main Realized Partisan Dilution asymmetry dose_party:initiative -0.00022 0.01365
12a Main Maximum Partisan Partisan bias dose_worst 0.00531 0.01739
12a Main Maximum Partisan Partisan bias nash_worst_10 -0.00102 0.00485
12a Main Maximum Partisan Partisan bias dvs_08 0.00932 0.03704
12a Main Maximum Partisan Partisan bias is_south 0.02323 0.01757
12a Main Maximum Partisan Partisan bias log(n_distr_20) 0.00575 0.01587
12a Main Maximum Partisan Partisan bias distr_chg 0.00257 0.01645
12a Main Maximum Partisan Partisan bias log(1 + corrupt) -0.01038 0.01185
12a Main Maximum Partisan Partisan bias initiative -0.00104 0.01372
12a Main Maximum Partisan Partisan bias drawer_ctrl_chg -0.00819 0.01749
12a Main Maximum Partisan Partisan bias court_ctrl_chg 0.00730 0.01346
12a Main Maximum Partisan Partisan bias dose_worst:nash_worst_10 0.00045 0.00214
12a Main Maximum Partisan Partisan bias dose_worst:dvs_08 -0.00069 0.01420
12a Main Maximum Partisan Partisan bias dose_worst:is_south 0.00693 0.01109
12a Main Maximum Partisan Partisan bias dose_worst:log(n_distr_20) 0.00089 0.00300
12a Main Maximum Partisan Partisan bias dose_worst:distr_chg -0.00089 0.00884
12a Main Maximum Partisan Partisan bias dose_worst:log(1 + corrupt) 0.00071 0.00229
12a Main Maximum Partisan Partisan bias dose_worst:initiative -0.00571 0.00961
12a Main Maximum Partisan Partisan bias dose_worst:drawer_ctrl_chg -0.00065 0.00752
12a Main Maximum Partisan Partisan bias dose_worst:court_ctrl_chg 0.00225 0.00888
12b Main Realized Partisan Partisan bias dose_party 0.00126 0.00705
12b Main Realized Partisan Partisan bias nash_party_10 0.00230 0.00343
12b Main Realized Partisan Partisan bias dvs_08 0.00140 0.03454
12b Main Realized Partisan Partisan bias is_south 0.00520 0.01869
12b Main Realized Partisan Partisan bias log(n_distr_20) 0.01052 0.01654
12b Main Realized Partisan Partisan bias distr_chg 0.00173 0.01667
12b Main Realized Partisan Partisan bias log(1 + corrupt) -0.01310 0.01174
12b Main Realized Partisan Partisan bias initiative -0.00087 0.01318
12b Main Realized Partisan Partisan bias dose_party:nash_party_10 0.00040 0.00092
12b Main Realized Partisan Partisan bias dose_party:dvs_08 0.00102 0.00770
12b Main Realized Partisan Partisan bias dose_party:is_south -0.00036 0.00382
12b Main Realized Partisan Partisan bias dose_party:log(n_distr_20) 0.00104 0.00218
12b Main Realized Partisan Partisan bias dose_party:distr_chg 0.00240 0.00425
12b Main Realized Partisan Partisan bias dose_party:log(1 + corrupt) 0.00057 0.00132
12b Main Realized Partisan Partisan bias dose_party:initiative 0.00177 0.00466
13a Main Maximum Partisan Differential partisan harm dose_worst 0.02297 0.01991
13a Main Maximum Partisan Differential partisan harm nash_worst_10 0.00497 0.00547
13a Main Maximum Partisan Differential partisan harm dvs_08 -0.03640 0.04470
13a Main Maximum Partisan Differential partisan harm is_south 0.01885 0.01986
13a Main Maximum Partisan Differential partisan harm log(n_distr_20) 0.01477 0.01666
13a Main Maximum Partisan Differential partisan harm distr_chg -0.02393 0.01787
13a Main Maximum Partisan Differential partisan harm log(1 + corrupt) -0.00693 0.01273
13a Main Maximum Partisan Differential partisan harm initiative 0.00572 0.01571
13a Main Maximum Partisan Differential partisan harm drawer_ctrl_chg 0.00395 0.01902
13a Main Maximum Partisan Differential partisan harm court_ctrl_chg -0.00338 0.01566
13a Main Maximum Partisan Differential partisan harm dose_worst:nash_worst_10 0.00060 0.00271
13a Main Maximum Partisan Differential partisan harm dose_worst:dvs_08 0.00155 0.01652
13a Main Maximum Partisan Differential partisan harm dose_worst:is_south 0.00453 0.01317
13a Main Maximum Partisan Differential partisan harm dose_worst:log(n_distr_20) -0.00008 0.00361
13a Main Maximum Partisan Differential partisan harm dose_worst:distr_chg -0.00443 0.01131
13a Main Maximum Partisan Differential partisan harm dose_worst:log(1 + corrupt) 0.00033 0.00279
13a Main Maximum Partisan Differential partisan harm dose_worst:initiative 0.00243 0.01099
13a Main Maximum Partisan Differential partisan harm dose_worst:drawer_ctrl_chg -0.00441 0.00946
13a Main Maximum Partisan Differential partisan harm dose_worst:court_ctrl_chg -0.00095 0.01050
13b Main Realized Partisan Differential partisan harm dose_party 0.00833 0.00797
13b Main Realized Partisan Differential partisan harm nash_party_10 0.00924 0.00375
13b Main Realized Partisan Differential partisan harm dvs_08 -0.02451 0.03799
13b Main Realized Partisan Differential partisan harm is_south 0.00017 0.01813
13b Main Realized Partisan Differential partisan harm log(n_distr_20) 0.01871 0.01804
13b Main Realized Partisan Differential partisan harm distr_chg -0.02270 0.01694
13b Main Realized Partisan Differential partisan harm log(1 + corrupt) -0.00943 0.01231
13b Main Realized Partisan Differential partisan harm initiative 0.00828 0.01460
13b Main Realized Partisan Differential partisan harm dose_party:nash_party_10 0.00096 0.00113
13b Main Realized Partisan Differential partisan harm dose_party:dvs_08 0.00172 0.00908
13b Main Realized Partisan Differential partisan harm dose_party:is_south -0.00096 0.00508
13b Main Realized Partisan Differential partisan harm dose_party:log(n_distr_20) 0.00178 0.00268
13b Main Realized Partisan Differential partisan harm dose_party:distr_chg 0.00279 0.00493
13b Main Realized Partisan Differential partisan harm dose_party:log(1 + corrupt) 0.00057 0.00159
13b Main Realized Partisan Differential partisan harm dose_party:initiative -0.00001 0.00525